### v3

# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

Abstract
We address the applicability of hybrid quantum-classical heuristics for practical railway rescheduling management problems. We build an integer linear model for the given problem and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison. The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments and covers all the requirements posed by the operator of the network. The computational results demonstrate the readiness for application and benefits of quantum-classical hybrid solvers in the realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation. At the same time, the solutions that were obtained were feasible. Moreover, though they are probabilistic (heuristics) they offer a valid alternative by returning a range of possible solutions the dispatcher can choose from. And, most importantly, they outperform classical solvers in some cases.
Keywords
[…]
1 Introduction
Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policy, leading to traffic challenges in passenger and cargo rail transport. The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which makes railway systems overloaded.
Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [23]. Examples of such disturbances include: late train departures/arrivals, extended dwell times, or (partial) track closures (also referred to as disruptions). These can last from several minutes up to hours. The impact of disturbances can propagate to multiple sections in the railway network (cf. the work of [50] and references therein). Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators. To limit these impacts as much as possible it is necessary to make proper rescheduling decisions quickly. The dispatchers need to reschedule and partially reroute the trains, aiming at the minimization of negative consequences. Still, in many places, dispatchers are using their own intuition or simple heuristics like FCFS (First Come First Served); resulting in decisions far from the objective of consequence minimization. In the last decades, there has been a growing research interest in mathematical optimization methods in support of rail dispatchers in decision making. While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [36], the maximal delay that cannot be avoided [14], or fuel consumption measures  [29].
As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for efficient suppression of the consequences of disturbances. Hence, large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time. Therefore it is vital to develop efficient algorithms. However, the current conventional rescheduling optimization models have difficulties addressing complex and large-scale instances in suitable computation time [8, 43, 35].
The railway dispatching/rescheduling problem can broadly be recognized as being equivalent to job-shop scheduling with blocking and no-wait constraints [49, 41]. A possible modeling approach is based on alternative graphs employing order and precedence variables [36], facilitating the formulation as a mixed integer linear program which is often large, hence, specialized algorithms are often used instead [35, 14]. An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants. Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling [8, 39, 43, 18].
As time indexed models result in integer or 0-1 programs, they are suitable [53] for to be solved on new types of hardware: quantum annealers [2, 33]. These physical devices can be considered as stochastic heuristic solvers for Quadratic Unconstrained Binary Optimization (QUBO) problems, and there are few commercially available options, including D-Wave [31]. The exploration of their potential applications attracts a growing research interest (c.f. Section 2.2), in which railway applications are not yet strongly represented.
In this paper, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple track railway lines with given planned train paths. Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections. We apply a new hybrid rescheduling algorithm combining classical-quantum modelling and based on Quantum Annealing (QA). This work extends on a particular linear modeling strategy, partly explored on a toy model in [21]. We develop an Integer Linear Programming (ILP) model, which is solved with proprietary D-Wave solvers as well as with CPLEX for comparison. The D-Wave approach is analysed in detail, with respect to its applicability and performance in a practically relevant situation. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready for tackling challenging railway rescheduling problems. During the preparation of the present work, another contribution has appeared in the literature [56], which is similar to our in that it sets up a QUBO model for a somewhat similar problem, and solves it by a simulation of (quantum) annealing. This contribution, however, considers a different railway infrastructure: high-speed trains, and considers timetable optimization, for which it employs a model different from the one described here.
The main contributions of the present work are the following:
  • A railway traffic dispatching model is introduced for rescheduling trains due to disturbances and disruptions taking into account given timetable, rolling stock connections and a network with single or multi-track segments.
  • QA-based hybrid heuristics are, for the first time, applied to real-life problems of railway rescheduling optimization on urban railway networks, these heuristics have some but still meaningfully input of real QA.
  • The performance of the hybrid (quantum classical) D-Wave solver is demonstrated on a real-life network in Poland and diverse disturbance scenarios of different sizes and types.
  • The D-Wave hybrid solver provides good quality solutions in short time, also for the complex instances in limited computational time.
  • For certain instances the hybrid (quantum classical) D-Wave solver outperforms classical solvers in terms of computational time, yielding still feasible solutions.
The paper is organized as follows. In Section 2 we review state-of-the-art literature and identify scientific gaps we intend to fill, in Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
2 Literature review
In this section, we present time-indexed models known from the literature, comparing them briefly with our modelling approach. For a general review of railway timetabling and rescheduling we refer to the reviews by [38] and [7]. Second, we review recent QA applications in optimization in general, and also in the rail/transport domain. Finally, we summarize the existing scientific gaps.
2.1 Classical railway dispatching/rescheduling
Time-indexed modelling of railway scheduling and rescheduling is quite common in literature for routing [39, 8] and scheduling [9, 29, 43, 13] . [8] focus on determining train’s path for a complex central railway station area (called the blocking - stairway), used a discrete-time model resulting in a $0-1$ program. Their model is successfully demonstrated on an operational day at the central railway station area Berne, Switzerland. Similarly, [39] consider train movements in a single major (and thus complex) railway junction, including a freight yard and a few minor stations. They build a mixed integer programming model that can also be solved close to optimal in case of practically relevant instances. They address the train movements on the level of detail of acceleration and deceleration strategies, hence adopting a significantly finer discrete time scale than our one-minute resolution. When compared to our problem, these two models addressed a high traffic in a complex station area. Instead, we are concerned more with a bigger area, multiple stations, and mixed track, but a lower train density and complexity. Our train routes fixed within the stations, and we also model shunting movements that is not a subject of the cited reference.
[9] proposed a graph theoretic formulation for the problem using a directed multigraph in which nodes correspond to departures/arrivals at a certain station at a given time instant. This formulation is used to derive an integer linear programming model that is solved using a Lagrangian relaxation. [13] introduced a new pure 0-1 programming formulation, and call it Tick Formulation, to model the Deadlock Detection (DD) problem. [43] focused “on the train dispatching problem on an N-track network, with the main challenge of how to formulate specific retiming, reordering, retracking and rerouting options in combination”. Their model was demonstrated on generic networks of different structure than what we address. In our situation we are limited to retiming and reordering of trains. Structurally, their model shows similarities to ours, including the role of binary variables and also the introduction of "cells" that generalize the objects the occupancy of which can be controlled: they can be one or more blocks, stations, etc. In our model, the decision variables will be explicitly linked to a subset of stations.
2.2 Quantum Annealing and its applications
Quantum Annealing (QA) [15] is an optimization method analogous to Simulated Annealing (SA). Both can be used for the unconstrained minimization of a quadratic function of ±$\pm 1$-valued variables, which represents the energy configuration for a set of spins of the Ising model [30, 4]; an important and celebrated problem in physics. The Ising problem can be equivalently rewritten in a form of a Quadratic Unconstrained Binary Optimization (QUBO) problem with binary ($0-1$) variables. The Max-Cut problem [5] is also equivalent to the Ising model or QUBO. The problem is NP-hard, has tremendous literature in classical optimization, and there is significant ongoing progress in the development of algorithms to solve it [22, 28]. An extensive list of Ising problems and their formulations are presented e.g. in the work of [37].
Hardware quantum annealers, like the D-Wave machine, implement a quantum version of the Ising model, assigning a quantum bit, i.e. a two-level physical quantum system to each bit. The system employs real physical two-level quantum systems with a tunable energy operator (Hamiltonian), the energy being the objective function of the optimization problem. An adiabatic evolution of the quantum system is implemented: the interactions between the spins are slowly changed from an initial Hamiltonian with a simple minimal-energy state to the Hamiltonian corresponding to the objective according to the rules of quantum mechanics. According to the adiabatic theorem of quantum mechanics [3], under certain conditions the physical system remains in its lowest energy state during the evolution, and an optimal, minimum-energy solution can be possibly read out at the end.
An ideal adiabatic quantum computer would be a device with a large number of perfect quantum bits completely decoupled from their environment, very close to zero temperature. Any pair of qubits could be coupled, and even, more general couplings, e.g. involving $3$ or more spins could take place. Such a system would exhibit perfect quantum coherence and entanglement. Even in such an ideal system the evolution time to reach at a minimum energy state with high enough probability is inverse proportional to the gap between the minimal energy configuration and the one closest to it. Certainly the gap is not known in advance, hence, the annealing time has to be determined by experimenting in practice. A more severe issue is that the gap can be very small , hence, it is not possible to solve all possible hard problems on such a setup. In spite of that, there are problems in which this approach can be efficient.
[…]
Operationally, from the optimizer’s point of view quantum annealers can be viewed as probabilistic heuristic solvers returning a statistical sample [20] of configurations which are supposed to be optimal or close to optimal. It must be stressed that the quantum annealers are not algorithms running on digital computers; they are analog devices implemented physically. This latter implies that the coefficients of the problem are encoded with a limited accuracy, and the algorithmic properties of the particular optimization problem such as its complexity class will not determine the solver’s behaviour.
To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently. Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ [11], including the one used in the present work, belong to this family. Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
Currently, the applicability of QA technologies is being explored. Benchmark problems are solved [55, 42], comparisons with digital computers are made [32], etc. The application of quantum annealers in transportation optimization is a new area with only a few contributions so far. There is an apparent interest in this research direction in the aviation domain [47, 51, 48], which relates also to the already mentioned benchmarking of annealers as done by[55]. Shipment rerouting was also addressed in the context of QA [58]. There are a few examples of hybrid solvers’ applications, too. These include traffic flow optimization [44], multi-car paint shop optimization [57], tail assignment problem [40], and vehicle optimization [24].
In railway context, the first applications of QA can be found in train dispatching/rescheduling [19, 21] and rolling stock planning [27]. In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [18]. A followup paper [21] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems. The present work continues this line of research.
2.3 Scientific gaps
We recognize several important scientific gaps. First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances. Second, the QA models for the studied optmimzation problem that have been introduced so far [19, 21] had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices. Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling.
In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks. To do so, we use a hybrid approach combining quantum and classical computing. For a fair comparison of current and future QA approaches with state-of-the-art classical approaches, we elaborate the railway model readily suitable for both classical and hybrid (quantum classical) solvers in a comparable manner. Quantum computing is believed to develop rapidly in near future. Hence, demonstration quantum readiness for railway problems right now is step forward towards the demonstration of quantum advantage in future. On future quantum devices the comparison with classical solvers is expected to be in strong favour for the quantum ones. In this way we also contribute in developing a novel general set of railway rescheduling models that has a potential to scale well with the size of the problem, aiming to overcome the curse of dimensionality.
3 Problem description
Figure 1: Edges and nodes on the example railway network. Interlocking areas of stations in green.
We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations. Also, we consider an urban railway network with mixed tracks from single, double, up to quadruple tracks.
We model a railway network with edges and nodes of a graph. Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple track lines). Each edge is composed of one or more tracks. Each track consists of blocks sections defined by pairs of signals [39]. Each block section can be occupied by at most one train at a time. A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains. We assume a 2-block signalling system, meaning that two free blocks are required between the consecutive trains. We also assume green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
Each node is composed of station tracks (blocks) and interlocking areas. One station track can be occupied by one train at a time. Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered in order to to guarantee that only one train from the pair can occupy the area at the time.
A train’s route is the sequence of blocks the train passes during its journey. A train path is a sequence of arrival and departure times of a particular train assigned to a train route. We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time or passing time, resulting in integer variables. This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning the inherently binary variables. The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to precedence of trains, encoded later on precedence binary variables.
In the network, two trains can follow each other, i.e. keep the given order, meet and pass (M-P) when going the opposite directions, and meet and overtake (M-O), i.e. change the order, when going the same direction. Single tracks are designed for bidirectional traffic, double-tracks for unidirectional, one for each direction, and multi-track lines can be combination of unidirectional and bidirectional. On single-track lines, M-Ps and M-Os are only possible at stations. To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single track line heading in opposite directions.
Trains in the same direction preserve their order between stations, and keep the minimal headway time between each other. To prescribe these we determine set $\mathcal{J}^{2}_{\text{headway}}$: the set of all pairs of trains that can potentially violate the minimal headway condition.
The train traffic is scheduled on the basis of the given timetable. The timetable contains all the train paths. In the train rescheduling problem, we are given an initial timetable that is conflict-free. However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc. Following [10], a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously. We assume that the conflict may occur either on the railway line or at the station. Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single track line by two trains heading in opposite directions at the same time. Conflicts at stations include claiming a station track by two trains at the same time, claiming a station switch (in the interlocking area) by two trains at the same time.
The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing and retiming for trains claiming for the same resource. Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable. These changes typically result in delays of additional trains that were not directly involved in the conflict otherwise, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable. It is common in the literature to minimize a function of these secondary delays (see e.g. the work of [14]); we will do this in the present contriubution, too).
An important element of our approach is that from amongst all stations we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only. Hence, as decision stations we select those stations, where routes of trains intersect, where trains start or terminate, or where selected part network is bounded. The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths has to be made with respect to these stations. (If we change some of the trains’ routes in a re-routing process, new model with new parameters if developed, e.g. additional decision stations may appear.) In what follows, by a station we always mean a decision station. This also means that non-decision stations appear only through the parameter values in the model, they do not appear as indices of decision variables or parameters. Headways, for instance, are calculated between decision stations, taking into account all the line blocks and station blocks of non-decision stations in between.
On the station where a train terminates or sets off, shunting is also modeled. The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa. The depot is treated as the station, and consider it as a black box, without detailed layout. We treat shunting movements as service train from depot to the starting station of the service train, or from the terminating station to the depot. The rolling stock circulation condition is applied to ensure the precedence between the service train and the actual train.
In this way our model shares features of both microscopic and macroscopic models. At the decision stations it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting. Meanwhile from the point of view of the rest of the network it is a macroscopic model.
4 Methods and Model
In the following we describe our model in detail. Section 4.1 defines sets, parameters and decision variables. Section 4.2 describes our Integer Linear Programming (ILP) formulation.
4.1 Sets, parameters and decisions
To formulate our decision variables, constraints and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan. Also, we introduce parameters calculated from the same input.
4.1.1 Sets
Let us denote by $\hat{S}$ the set of all stations, and by $S\subset\hat{S}$ the decision stations we have determined in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. Set $\mathcal{S}_{j}$ includes decision stations only and it is an ordered set. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following. Set $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J}\left(\forall s\in\mathcal{S}\right)$ is the set of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair. This set is deduced from the rolling stock circulation plan and the timetable. Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains which are close enough to each other in time so that precedence variables have to be defined for them. More details will be given at the description of the parameter $d_{\max}$ which this set depends on. Set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that can potentially violate the minimal headway time between two subsequent trains on a line segment. For a pair of trains to be in $\mathcal{J}^{2\,(\text{headway})}$, both their routes need to include the same line segment so that the trains are moving in the same direction on it and can meet there according to model parameters, i.e. maximal allowed delay. Set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of all trains that share at least one single-track line segment as a common part of their routes so that they are heading in the opposite direction. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to occupy the same track on station $s$ anytime during the planning horizon, thereby competing for the same track. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their departure from station $s$. Set $\left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ in order to have $j$ to departure $s$ while $j^{\prime}$ arrive $s$ from the direction of $s^{\prime}$. Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-P possibility for them between $s$ and $s^{\prime}$ Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ so that either they both come from the direction of $s^{\prime}$ but there is a M-P possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. Set $\left(\forall j\in\mathcal{J}\right)\ \mathcal{C}^{2}_{j}$ is the set of all station pairs that are subsequent in the route of $j$. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. All these sets can be enumerated on the basis of the input data in a straightforward manner.
4.1.2 Parameters
The parameters that appear in our model are the following:
  • $\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ is the running time of $j$ from $s$ to $s^{\prime}$.
  • $\tau^{\text{(headway)}}(j,j^{\prime},s\rightarrow s^{\prime})$ is the minimal headway time for $j^{\prime}$ following $j$ from $s$ to $s^{\prime}$.
  • $\tau^{(\text{switch})}(j,j^{\prime},s)$ is the running time of $j$ over a interlocking area of $s$, where $j$ may be in conflict with $j^{\prime}$. (In our examples we do not consider cases when, e.g., on bigger stations there are multiple switches with different technological times. The model could trivially be extended to cover such scenarios by adding extra indices if relevant.)
  • $\tau^{(\text{dwell})}(s,j)$ is the minimal dwell time of $j$ at $s$.
  • $\tau^{\text{(turn)}}(s,j,j^{\prime})$ is the minimum turnaround time for the rolling stock of a train $j$ terminating at $s$ to continue its journey as train $j^{\prime}$.
  • $\sigma(j,s)$ is the scheduled departure time of $s$ from $j$, given by the original timetable.
  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance and/or disruption in the network happens. It is the maximum of $\sigma(j,s)$ and the technically feasible earliest departure time, and may be more constraining than the first depending on the initial delays. The latter is calculated assuming that the train follows its planned route at minimum running time, and that there are no other trains present on the network.
  • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$. This parameter sets an upper limit for the possible secondary delays that can arise on the particular part of the network. Setting such a bound is common in the literature, see e.g. the work of [14]. We set the same value of this parameter for all our testing instances. It has to be big enough so that no secondary delay exceeds it (e.g. $d_{\max}=40$ excludes the possibility of a one-hour delay due to waiting for another delayed train). Meanwhile it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $J^{2\,(\text{close})}$. A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed/disrupted timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$. The number of constraints will be proportional to the average number of trains that can meet another train at a station; this will be linear in $d_{\max}$.
4.1.3 Decision variables
Our decision variables are the following. First, we use departure time variables:
$t^{(\text{out})}(j,s)\in\mathbb{N}$ (1)
defining the departure time of train $j\in\mathcal{J}$ from station $s\in\mathcal{S}_{j}$. Such a variable is defined for all decision stations $\mathcal{S}_{j}$. In the case of trains that terminate within the modeled part of the network, the last station is not included in $\mathcal{S}_{j}$ The arrival time of the trains at stations, $t^{(\text{in})}(j,s)\in\mathbb{N}$ are trivially related to the $t^{(\text{out})}(j,s)$ variables through a constant offset (c.f. Eq. (3)), given the fixed running time assumption.
Second, in addition to the time variables, we use three sets of binary precedence variables. Decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: the variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and 0 otherwise. The $(j,j^{\prime},s)$ tuples for which an $y$ variable is defined will be specified later. Similarly, the precedence variables $y^{(\text{in})}(j,j^{\prime},s)\in\{0,1\}$ prescribe the order at the entry to stations; the value is $1$ if $j$ arrives to $s$ before $j^{\prime}$. The last set of precedence variables describes the precedence of trains at some resource located between stations $s$ and $s^{\prime}$ (e.g. single track used by $j$ and $j^{\prime}$ heading in opposite direction). The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$. Also in this case, the quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later.
4.2 ILP formulation
Given the index sets, variables, and the parameters and decision variables of the model, now we formulate the optimization objective and the constraints.
4.2.1 Objective function
Our goal is to minimize the secondary delays that occur in the analyzed part of the railway network. Hence, a suitable objective function is the weighted sum of secondary delays at the destination station:
$f(t)=\frac{1}{d_{\text{max}}}\sum_{j\in\mathcal{J}}w(j)\left(t^{(\text{out})}(j,s^{*})-\upsilon(j,s^{*})\right).$ (2)
where $s^{*}$ is the last element of $S_{j}$, and $w_{j}$ are weights for each train representing its priority. The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as the weighted sum of secondary delays on all stations is straightforward. The constant multiplier $1/d_{\text{max}}$ is optional; we use it for better comparability of different instances.
4.2.2 Constraints
The constraints of the model are the following.
Minimal running time
Each train needs a minimal time to get to the subsequent station:
$\forall j\in\mathcal{J}\ \forall(s,s^{\prime})\in\mathcal{C}_{j}\quad t^{(\text{in})}(j,s^{\prime})=t^{(\text{out})}(j,s)+\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ (3)
Headways
A minimal headway time is required between subsequent train pairs on the common part of their route as
$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}t^{(\text{out})}(j^{\prime},s)\geq$
$t^{(\text{out})}(j,s)+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})-C\cdot y^{(\text{out})}(j^{\prime},j,s),$ (4)
where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$. (Constraints of this structure are often termed as big-M constraints in this context.) In our implementation we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
$C=-\min(t^{(\text{out})}(j^{\prime},s))+\max(t^{(\text{out})}(j,s))+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})$ (5)
Single-track occupancy
Trains moving in opposite directions cannot meet on the same single-track line segment:
$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{single})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$
$t^{(\text{out})}(j^{\prime},s^{\prime})\geq t^{\text{(in)}}(j,s^{\prime})-C\cdot z(j^{\prime},j,s^{\prime},s),$ (6)
where $C$ is a big enough constant chosen similarly to that in Eq. (4).
Minimal dwell time
Each train has to occupy the station node for a prescribed time duration at each station:
$\forall j\in\mathcal{J}\forall s\in S_{j}\quad t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s).\$ (7)
Timetable
No train is allowed to depart before its scheduled departure time:
$\forall j\in\mathcal{J}\forall s\in S_{j}\quad t^{(\text{out})}(j,s)\geq\sigma(j,s).$ (8)
Station track occupancy
Station tracks can be occupied by at most one train at a time:
$\small{\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
where $C$ is chosen similarly to that in Eq. (4) again. Note that this requirement may not be needed for depot tracks; the exceptions can be handled by the proper definition of $\mathcal{J}^{2\,(\text{track})}_{s}$.
Interlocking area occupancy
These ensure that trains cannot meet in interlocking areas:
$\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}t^{(\text{out})}(j^{\prime},s)\geq$
$t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$ (10)
$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
$t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot z(j^{\prime},j,s^{\prime},s),$ (11)
$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
$t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s^{\prime}),$ (12)
$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
$t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{in})}(j^{\prime},j,s),$ (13)
with a choice of $C$ similar again to that in Eq. (4).
Rolling stock circulation constraints
These are introduced in order to bind the train with the shunting movement called also service train. If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
$\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s).$ (14)
Order of trains
We have additional conditions on the $y$-variables, concerning the case when M-O is not possible on the line or station.
$\begin{split}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}\\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\\ y^{(out)}(j,j^{\prime},s)=y^{(out)}(j,j^{\prime},s^{\prime})\\ \forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\\ y^{(in)}(j,j^{\prime},s^{\prime})=y^{(out)}(j,j^{\prime},s^{\prime})\end{split}$ (15)
The objective function in Eq. (2), together with the constraints in Eq. (3)-(14) define our mathematical programming model.
The model yields an integer linear program, and the time variables can be constrained even into a finite range using the parameter $d_{\max}$. Note that the binary variables have the obvious symmetry property
$\begin{split}y^{(\text{out})}(j^{\prime},j,s)=1-y^{(\text{out})}(j,j^{\prime},s)\\ z(j^{\prime},j,s^{\prime},s)=1-z(j,j^{\prime},s,s^{\prime}),\end{split}$ (16)
which is taken into account explicitly by using independent binary variables only.
As for the scaling of our model, the number of time variables i.e. #$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  [21]). In order to obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations. We assume that in average each train visits $\alpha\in(0,1)$ fraction of the stations, we have then
#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}.$ (17)
as a tighter bound.
Suppose, that most of the network is composed mainly of double track lines. Then there will be typically a single precedence variable per train and station ($y^{(out)}(j,j^{\prime},s)$), and thus their number can be estimated by:
#$\#y+\#z\approx\#y\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S},$ (18)
where $N(d_{\max})$ is number of trains the train can meet at a station on average. Under such assumptions the total number of variables grows linearly with number of trains and stations:
#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}\#\mathcal{S}.$ (19)
If, on the other hand, the network has dominantly single track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
#$\#y=\#z\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S}.$ (20)
In both cases we assume that there are not many ($y^{(in)}(j,j^{\prime},s)$) variables, as they are tied certain particular interlocking area conditions.
Based on Eq. (16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single track occupation constraints. For the interlocking area occupation constraints it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable. For the running time and minimal dwell time we expect one constraint per train and station. Thus number of constraints can be estimated by
#$\#constr.\approx 6\#y+2\#z+2\#\mathcal{J}\#\mathcal{S}$ (21)
where the second term is tied to single track line conditions (one condition per $z$ variable).
#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single track}.\end{cases}$ (22)
Because of the limitations in time imposed by $d_{\max}$ the model can be viewed as linear in number of trains and number of stations. Hence, as expected, the model is more complex for the network in which single track lines dominate. Although the analysed network is mostly composed of the double track lines, we will analyse certain use-cases with the fraction of single track lines increased.
Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables and number of constraints in the number of trains.
5 Hybrid quantum-based approach
To overcome the limitations described in in Section II of current hardware quantum annealers, we apply hybrid (quantum classical) solver. In particular, we use the ’Leap Hybrid Solver Service (HSS)’ [11]; a cloud-based proprietary solver which is developed by the market leader of QA hardware and is available as a service. It utilizes a hybrid approach that combines classical computational power with quantum processing.
In particular, we have used the Constrained Quadratic Model (CQM) [12] Solver. The CQM hybrid solver works as follows.
Input
A constrained linear or quadratic model, and the value of the solver parameter t_min, which prescribes the minimum required run time (in seconds) solver is allowed to work on any problem.
Preprocessing
Preprocessing includes subproblem identification and decomposition to smaller instances [52]. Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
Processing
The hybrid solver implements a workflow that combines a portfolio of various classical heuristics (including tabu search, simulated annealing, etc.) running on powerful CPUs (Central Processing Units) and GPUs (Graphics Processing Units). In the course of the solution process the hardware quantum annealer is also invoked, in order to solve smaller QUBO problems that can potentially boost the classical heuristics. Integer problems are transformed to binary using state-of-the-art methods like the one described by [25][34], generally applicable for integer problems. When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions. After these readouts, the sample is incorporated into the solution workflow. In this way, even though the physical annealer supports problems with limited size (e.g. the Pegasus machine has $5500$ physical qubits [16], and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process. The final solution is composed by the solver’s module.
Output
The solution of the integer or quadratic constrained problem. The workflow results in a solution that is not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver. Furthermore, the solver is probabilistic, hence each at each run the final solution may be different.
While D-Wave’s proprietary solvers act as black boxes, i.e. they hide the exact details of the described process process from the user, the output is supplemented with timing parameters, including run time: the total elapsed time including system overhead, and Quantum Processing Unit (QPU) access time which is is the time spent accessing the actual quantum hardware. These parameters enable a comparison with other solvers.
As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min. This is a time limit for heuristics running in parallel[12]: unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution. We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting to the model performance and solution quality for our problem.
The CQM Solver can solve problems encoded in the form of Eq. (2) – (15) (c.f. Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints. Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum classical) solvers. As a classical solver, IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) was used. The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory. As hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.
6 Computational results
In this section, we demonstrate the performance of the proposed hybrid (quantum classical) approach on a network at Polish railways - central part of the Upper Silesia Metropolis. In particular, we compare the performance of DWave’s hybrid quantum-classical solver (CQM) with CPLEX. The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances and disruptions including various initial delays and track closures. As the objective is tied to secondary delays, we can roughly apply it to asses the degree of difficulty of the problem.
Section 6.1 describes the considered network and traffic characteristics. Section 6.2 presents results concerning generic examples on the core line of the network, comparing double track line scenarios with single track line scenarios and closure scenarios. Section 6.3 presents analogous results for larger network with real-life timetable. Here the case studies concerns various scenarios with different degree of difficulty.
6.1 Railway network
We use the Open Railway Map [^1] https://www.openrailwaymap.org/, visited 2022-11-25 representation of the selected part of the Polish railway network located in the central part of the Metropolis GZM (Poland), presented in Fig 2. The selected part of the network combines heterogeneous network layout including single-, double-, triple-, and quadruple-track lines. It comprises 25 nodes including $11$ stations and $3$ junctions, $146$ blocks, and 2 depots. The infrastructure is managed by Polish state infrastructure manager PKP PLK (Polskie Koleje Państwowe, Polskie Linie Kolejowa eng. Polish State Railways, Polish Railway Lines).
Figure 2: Selected part of Polish railway network located in the central part of the Metropolis and examined in our research.
The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analysed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O). As described before, the decision stations are the only ones that appear explicitly in our model. Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made with respect to the decision stations. All other stations and junctions are treated as line blocks as long as no rerouting or retracking is considered; they will appear in the model implicitly via parameters. These stations include:
[…]
In all computation priority weights in Eq. (2) we use $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable). As for the solver parameter t_min, after initial experiments we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality. Sensitivity analysis regarding this parameter is performed in Appendix A for selected instances.
6.2 Synthetic experiments on the selected railway line
As the first set of experiments we address, synthetic instances on a part of the network - railway line KO-GLC, see Fig 2. The goal in these scenarios is to compare the CPLEX and CQM solvers for: the double track line ($line1$), the double track line with closures ($line2$), and the single track line ($line3$). Shunting movements and rolling stock circulation are not considered in this set of experiments. In particular, the addressed configurations are the following:
  1. 1.
    $Line1$, the double track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
  2. 2.
    $Line2$, similar to $line1$, with the additional closure of one of the tracks between ZZ and RCB. We use a cyclic $2$-hour timetable with $10$ trains each hour and each direction, i.e. we have $40$ trains.
  3. 3.
    $Line3$, the whole line is considered as a single track, with a timetable of $3$ hours traffic and $21$ trains.
For $line1$ and $line3$, the original timetable is feasible, while for $line2$ it is not feasible due to the closure. For each of the lines we compute $12$ instances, each with different initial delays of trains subsets. Instance 00 yields no initial delays. Thereafter, we use the parameter value $d_{\max}=40\ \text{min}$, in all cases in this and subsequent subsection. Assuming that each train interacts with $N(d_{\max})=12$ trains for $line1$ and $line2$, and with $N(d_{\max})=6$ trains for $line3$, Eqs. (17) - (19), (22) allows us to quickly estimate the sizes of the problem. This sizes are compared with computed (comp.) actual number of variables and constraints (mean over instances) in Tab. 1. Here, $line2$ is the most complex in terms of constraints, what is reflected later by computational difficulties of some of its instances. Furthermore, $line2$ is partially double track and partially single track, hence the actual number of constraints is bounded from above by the estimate for the single track line and from below by the estimate for the double track line. $Line1$ includes double-, whereas $line3$ single-track segments only. There actual number of constraints and variables appeared to be close to the estimate.
l #$\#\mathcal{S}$$\#\mathcal{J}$ int vars # precedence vars # constraints
comp. mean Eq. (17) act. mean
Eq. (18)/Eq. (20)
doub./sing.
comp. mean
Eq. (22)
doub./sing.
1 3 59 118 118 1538 1415 / - 7889 8731 / -
2 5 40 142 133 1394 1600/3200 8491 6867/13067
3 5 21 79 70 592 - / 840 3057 - / 3499
Table 1: Synthetic experiments. The first column is the ordinal number of the line. For the estimated and actual number of variables and constraints, we use $\alpha=2/3$ in Eqs. (17) - (19), (22) (doub. - double track, sing. - single track, comp. mean - computed actual number of variables or constraints, mean over instances).
The computational results are presented in Figs. 3,  4, and 5 for $line1$, $line2$, and $line3$, respectively. Each figure shows objective $\times$ $d_{\text{max}}$ (top part), computation time (comp. time, middle part) and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s. CPLEX results are also displayed.
For scenario $line1$ (Fig. 3), CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$). Similarly, for $line3$ (Fig. 5), only small deviations from the optimum can be observed for a number of instances. All computational times for $line1$ and $line3$ were significantly below $0.4$min, although CPLEX performed faster. Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objectives but t_min$=20s$ requires almost $4$ times longer computational time.
For $line2$ (Fig. 4), all CQM solutions were suboptimal, although feasible. More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than CPLEX, in several cases up to 5×5$\times$5 × faster. This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5). This observation is approved by more detailed discussion on t_min parameter in Appendix A. In principle, $line1$ (Fig. 3) and $line3$ (Fig. 5) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster. We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems.
Figure 3: Synthetic experiments $Line1$, comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Total computational time (middle panel) and QPU times (lower panel) were provided by the D-Wave output. Results of both solvers were similar in terms of the objective, but the CPLEX computation was faster.
Figure 4: Synthetic experiments $Line2$, comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. Note that the CQM is fast, even when compared to CPLEX. The found solutions are suboptimal but can be useful. All the displayed solutions are feasible. $Line2$, namely the double track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective). On this example, there are instances, where the current CQM D-Wave hybrid solver outperforms CPLEX in terms of computational time. Increasing the value of the t_min parameter in most cases improves the quality of the solution at the cost of computational time.
Figure 5: Synthetic experiments $Line3$, comparison of the performance of classical solvers (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
6.3 Real-life experiments on the considered rail network
In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on presented network. We will also focus on the track closure situation, turning a double track line segment into a single track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2. Our computations address various use cases of train delays outside the network and within the network as well as the partial track closure. We constructed $9$ networks with increased level of difficulty ranging from small delays to disruptions, i.e. it is expected more trains are involved and/or disturbances are spread more over the network. These networks are tabulated in Table  2. Their details are as follows:
  1. 1.
    Networks $1$ $3$ concern only disturbances due to delayed trains and no closures.
  2. 2.
    Networks with number $4$ or higher concern also disturbances due to closures.
    1. (a)
      Networks $4$ and $5$ concerns rerouting trains from double track line KTC-CB to the single track line with higher passing times, see Fig. 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
    2. (b)
      Network $6$ concerns multiple closures, i.e. change of multiple line KZ-KO(STM) and double track line KO-KL to single track lines. (See Fig. 2; the reason can be e.g. upcoming reconstruction works on this part of network.)
    3. (c)
      Networks $7$$9$ concerns closures of both $5$ and $6$, but each with different initial delays of trains. The intention of the last $3$ networks is to create a really challenging rescheduling problem.
  3. 3.
    For comparison, network 00 is the default problem with no disturbances.
network #$\#S$ place trains involved description
A. Disturbtions from outside the network
1 10 KO-Ty IC 14006 KS 94113 Delayed IC (higher priority) is in conflict with on-time KS
2 10 arr. Ty KS 94766 KS 40518 IC 41004 KS 44862 IC 4120 Delayed 5 trains on the arrival to Ty, random delays of several minutes
B. Disturbtions originated within the network
3 10 whole network 14 trains KS + IC Random delays of 14 trains delays up to 35 minutes
C. Disturbtions within the network and closures
4 11 KTC-CB 10 trains KS + IC Double track line KTC-CB closed trains rerouted to single track line
5 11 whole network most trains Double track line KTC-CB closed as in 4 + delays as in 3
6 10 whole network all trains Multiple track KZ - KO(STM) and double track KO - KL changed into single track + delays as in 3
7 11 whole network all trains Multiple track KZ - KO(STM) and double track KO - KL changed into single track + double track line KTC-CB closed as in 4 + delays as in 3
8 11 whole network all trains Closures as in 7 + random delays of 13 trains up to 30 minutes
9 11 whole network all trains Closures as in 7 + random delays of 15 trains up to 30 minutes
Table 2: Real-life experiments. Networks of particular dispatching problems. In networks 00 - $5$ most of analysed network is double tracked, from network $6$ upward the number of single track lines in the network increases at the expense of the number of double and multiple track lines. For each network we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains. Beside the #$\#S$ stations there are $2$ depots in each network.
n. CPLEX
CQM hyb. t_min = $5$ s
mean value over $5$ realis.
comparison
CQM vs. CPLEX
#vars /
# constr.
obj. $\times$
dmax
comp.
time [s]
obj. $\times$
dmax
comp.
time [s]
QPU
time [s]
obj.
diff. %
comp t.
diff %
0 556 /1756 0.0 0.07 0.0 5.19 0.03 0 % -
1 556 / 1756 1.0 0.08 1.0 5.24 0.02 0 % -
2 556 / 1740 6.0 0.08 6.0 5.37 0.03 0 % -
3 556 / 1769 7.5 0.09 7.5 5.05 0.03 0 % -
4 662 / 2210 78.25 0.20 82.70 5.10 0.03 -5.6 % -
5 662 / 2204 114.75 0.25 132.55 5.25 0.02 -15.5 % -
6 711 / 2599 91.25 0.41 142.3 5.14 0.03 -55.9 % -
7 817 / 3029 188.75 7.98 263.4 5.12 0.02 -39.5 % 35.8 %
8 817 / 3074 157.75 3.70 271.65 5.12 0.02 -72.2 % -38.4 %
9 817 / 3081 185.5 6.51 263.85 5.11 0.02 -42.2 % 21.5 %
Table 3: Real-life experiments. Results of ILP optimization on CPLEX and CQM D-Wave hybrid solver. $d_{\max}=40$ was set for all networks. Computational times and QPU time were reported in D-Wave’s output. We have selected t_min = $5$s as it provides the best balance between computational time and the quality of solution. For detailed analysis see Appendix A. This appendix shows a detailed analysis of the impact of t_min on computation time and solution quality for two specific real-life lines $7$ and $9$. The last columns present a percentage comparison (CQM - CPLEX ) / CPLEX, hence, the positive values represent CQM advantage, whereas negative values show CPLEX advantage.
Table 3 tabulates the results comparing CPLEX and CQM in terms of objective function value and computation time; problem sizes and QPU times are included. As CQM is probabilistic, $5$ runs have been performed for each network, yielding $5$ different realizations. This also facilitates the analysis of the statistics of the output and the differences between particular realizations.
The level of difficulty increases with respect to the number of variables and constraints. Observe that CQM always returns a feasible solution. For the networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX. As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$). Meanwhile those of CQM remain almost constant around $5.1s$. QPU times remain not greater than $0.032s$. Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver even outperforms CPLEX with respect to computational time. It does not find the actual optimum, just a feasible solution close to it. For example, in network $9$, the computed CQM solution is around 30% worse than the optimal one. In conclusion, we expect benefits from the application of the hybrid solver on medium scale railway networks with multiple closures, as in networks 7-9. This is in accordance with our observation on synthetic data in Fig. 4, where the benefits appeared for more complex instances, too.
To evaluate the extent of difficulty of the problems analysed in Tab. 3, Tab. 4 presents mean (over stations and realizations) delays, resulting from CQM solutions, on selected stations (i.e. stations in central part of the network in Fig. 2). Meaningful delays are introduced starting from network $4$ onward, i.e. for networks with closures in a limited number of stations, see also Tab. 2. From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Tab. 2. However, for networks $7$ $9$ the mean delays are the highest and almost uniformly distributed among all stations, showing that the spread of disruption impacts the whole network. These have appeared to be more challenging network disruptions.
Fig. 6 shows more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (top panel) vs CQM (bottom panel). For network $4$ (left panel of Fig. 6), bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller. This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Tab. 2), and thus the spread is limited to these 3 stations. For network $7$ (right panel of Fig. 6), delays are more spread through the network due to disruption and initial disturbances. To conclude, we can expect advantage from a hybrid (quantum classical) approach particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network.
Delay [min]
network KO(STM) KO CB KL MJ
0 3.7 1.0 1.7 1.5 0.0
1 3.1 1.3 2.1 1.5 0.0
2 3.8 1.3 2.0 1.3 2.0
3 2.6 2.1 1.4 0.7 1.5
4 6.6 4.4 7.4 1.6 0.0
5 7.1 6.2 10.7 1.0 2.7
6 6.7 6.7 3.2 8.7 4.1
7 9.8 10.7 12.5 9.5 9.9
8 10.4 10.5 10.0 10.5 7.6
9 10.0 10.3 11.5 9.8 9.6
Table 4: Real-life experiments. Mean (over trains and realizations) secondary delay resulted from CQM solutions (Tab. 3) for all networks in Tab. 2. Non-zero values for network 00 comes from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes delay but make it up and leaves the network on time - these are typical issues for the railway traffic on analysed network.
Figure 6: Real-life experiments. Statistics of CPLEX (top panel) and CQM (bottom panel) solutions (Tab. 3) for selected networks in Tab. 2. The statistics were calculated over trains in the classical cases, and over trains and realizations in the CQM cases. Vertical lines are ranges, while horizontal lines are $0.25$ and $0.75$ percentiles.
In order to further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Figs. 7,  8, and 9, respectively. We have chosen the line segment between GLC and KZ, offering a good demonstration. (We have introduced a small artificial ’distance’ between KO and KO(STM) wich are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution. All the three proposed solutions have similar figures of merit, they all provide a valid option to be realized. The conflicts are resolved in somewhat different ways, see e.g. the paths of trains 40150, 5312, and 26103. In details:
  • in the CPLEX solution, train 26103 passes first the KTC-CB closure heading for GLC, then 40150 followed by 5312 go in opposite directions,
  • in the CQM solution, first realization, 40150 and 5312 pass first the KTC-CB closure and 26103 waits at KTC for them to pass,
  • in the CQM solution, second realization, 26103 passes first KTC-CB closure, then 5312 followed by 40150 go in opposite directions.
This results in following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$. Although, the first solution is optimal, it is reasonable to supply dispatchers with the range feasible solutions. Hence, these diagrams proved that CQM hybrid solver can produce usable solutions for railway traffic management.
An additional benefit of the probabilistic nature of the quantum-based method is that it produces different alternatives after each realization run without any modification. These could be offered to the dispatcher in a decision support system as possibilities. The dispatcher’s choice can be influenced by other circumstances not covered by the present model.
Figure 7: Real-life experiment. Time-distance diagram CPLEX solution of network $7$, see Tab. 2.
Figure 8: Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) first realization. Closures (yielding single track line) are between CB and KTC as well as between KO(STM) and KZ.
Figure 9: Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) second realization.
7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular we have encoded rescheduling problems in a linear integer program, and solved these using a state-of-the-art solver (CPLEX), as well as on the QA based CQM hybrid (quantum classical) solver. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better. This supports the expectation that future quantum devices will be efficient in solving large scale-problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that are beyond the scope of current exact models and heuristics.
While the quantum-based solvers would possibly return suboptimal solutions, these are still often better than the solutions obtained manually or based on smaller scale models. More importantly, all generated solutions were feasible. In each computation, the CQM hybrid solver always has taken some advantage of the QPU. Hence, we believe that such a little help from the QPU must have boosted the classical heuristics in the hybrid solver.
Recall that quantum annealer devices are subjects of many ongoing debates. Previous studies claimed that the quantum nature of their operation is limited [45], while recent works argue that their operation relies more on the thermalization with a cold environment in place of actual quantum adiabatic evolution [6]. They are often criticized for the small size of problems they can address and various classical solvers do outperform them on certain problems at the present state of the art. In spite of all these, our computational results demonstrate the readiness of hybrid (quantum classical) solvers to handle real-life railway problems. We believe that future investigations shall focus on hybrid approaches both for railway applications as well as more generally.
Finally, several future research directions can be determined. First concerns the ongoing worldwide works on the development of fault tolerant quantum computing devices [17] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular). Second concerns creation of custom open source hybrid solver that is dedicated for railway problems and will use noise of quantum device to our advantage e.g. by modelling some stochastic behaviour on railways. Third concerns evolution and application of quantum inspired techniques such as tensor networks technique, simulated bifurcation [26]. With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.
Data availability
The code and the data used for generating the numerical results can be found in https://github.com/iitis/railways_dispatching_silesia under DOI 10.5281/zenodo.10657323
Acknowledgements
This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no. BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK). This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD). For the purpose of Open Access, the authors has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
We would like to thank to Sebastian Deffner as well as to the whole scientific community of the Department of Physics, University of Maryland, Baltimore County (UMBC) for valuable discussion; and to Bartłomiej Gardas, and Zbigniew Puchała for valuable motivation; and to Katarzyna Gawlak, Akash Kundu, and Özlem Salehi for data validation and assistance with coding. We acknowledge the cooperation with Koleje Ślaskie sp. z o.o. (eng. Silesian Railways).
Compliance with ethical standards
Conflict of interest
The authors declare that they have no conflict of interest.
Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Appendix A. Comparative analyses for different values of the t_min parameter on selected solutions
The right choice of the t_min parameter can improve results meaningfully. Hence, in this Appendix we analyse in more detail the role of the t_min solver parameter for selected problems in Section 6.3. In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig. 10 and Fig. 11. The objective value is roughly log-linear in computational time. For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially. Overall, the improvement in the objective is overwhelmed by the increase of the computational time. This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling. Such a behaviour is plausible in the case of QA methods [46, 1, 20].
Figure 10: Real-life experiments. The sweep over t_min parameter for network $7$, see Tab. 2. All solutions were feasible. For higher t_min we expect smaller objective values but higher total computational time. For each parameter value there were $5$ realizations of the experiment were performed. Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
Figure 11: Real-life experiments. The sweep over t_min parameter for network $9$, see Tab. 2. All solutions were feasible, results were similar to those in Fig. 10.

### v4

@@
 Keywords
-[…]
+railway rescheduling, conflict management, heterogeneous urban railway network, quantum annealing, hybrid quantum-classical heuristics
 1 Introduction
 Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policy, leading to traffic challenges in passenger and cargo rail transport.
-The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which makes railway systems overloaded.
-Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [23].
-Examples of such disturbances include: late train departures/arrivals, extended dwell times, or (partial) track closures (also referred to as disruptions).
+The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems.
+Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays (Ge et al., 2022).
+Examples of such disturbances include: late train departures/arrivals, extended dwell times.
 These can last from several minutes up to hours.
-The impact of disturbances can propagate to multiple sections in the railway network (cf.
-the work of [50] and references therein).
+A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too (Cacchiani et al., 2014).
+In the present work we focus on rescheduling of trains, which is the network operator’s perspective, even though we also consider (partial) track closures, which are are typically considered as disruptions.
+In our examples, however they will be managed by modifying existing train paths.
+The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f.
+Törnquist (2007) and references therein).
 Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
-To limit these impacts as much as possible it is necessary to make proper rescheduling decisions quickly.
-The dispatchers need to reschedule and partially reroute the trains, aiming at the minimization of negative consequences.
-Still, in many places, dispatchers are using their own intuition or simple heuristics like FCFS (First Come First Served); resulting in decisions far from the objective of consequence minimization.
-In the last decades, there has been a growing research interest in mathematical optimization methods in support of rail dispatchers in decision making.
-While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [36], the maximal delay that cannot be avoided [14], or fuel consumption measures  [29].
 As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for efficient suppression of the consequences of disturbances.
-Hence, large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time.
-Therefore it is vital to develop efficient algorithms.
-However, the current conventional rescheduling optimization models have difficulties addressing complex and large-scale instances in suitable computation time [8, 43, 35].
-The railway dispatching/rescheduling problem can broadly be recognized as being equivalent to job-shop scheduling with blocking and no-wait constraints [49, 41].
-A possible modeling approach is based on alternative graphs employing order and precedence variables [36], facilitating the formulation as a mixed integer linear program which is often large, hence, specialized algorithms are often used instead [35, 14].
+While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays (Lange and Werner, 2018), the maximal delay that cannot be avoided (D’Ariano et al., 2007), or fuel consumption measures  (Harrod, 2011).
+Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time.
+The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints (Szpigel, 1973; Mascis and Pacciarelli, 2002).
+It has an extensive literature, already summarized in a number of reviews, including those by Cacchiani et al.
+(2014), by Lamorgese et al.
+(2018), or by Philip and Swapnesh (2022).
+The maybe most frequently applied modeling strategy is based on alternative graphs (D’Ariano et al., 2007; Mascis and Pacciarelli, 2002) which can be formulated as as a mixed integer linear program, often referred to as the big-M model.
+As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and the combinations of these; see the work of Lamorgese et al.
+(2018) for a review of the model and the methods.
 An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants.
-Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling [8, 39, 43, 18].
-As time indexed models result in integer or 0-1 programs, they are suitable [53] for to be solved on new types of hardware: quantum annealers [2, 33].
-These physical devices can be considered as stochastic heuristic solvers for Quadratic Unconstrained Binary Optimization (QUBO) problems, and there are few commercially available options, including D-Wave [31].
-The exploration of their potential applications attracts a growing research interest (c.f.
-Section 2.2), in which railway applications are not yet strongly represented.
-In this paper, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple track railway lines with given planned train paths.
+Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling (Caimi et al., 2012; Lusby et al., 2013; Meng and Zhou, 2014).
+As this approach results in a $0-1$ program, which is promising when one considers the use of $0-1$ solvers such as quantum hardware.
+Indeed, job-shop problems addressed on pure quantum hardware (Venturelli et al., 2015), and even in certain hybrid approaches (Kurowski et al., 2020) use similar models, and our previous work (Domino et al., 2023) which was the first to apply a quantum solver to railway dispatching problems also followed this route.
+In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP).
+We introduce a variant of this type of model which maintains a reasonable model size while addressing practically relevant problem instances.
+In this way, a hybrid classical-quantum integer solver can be applied to solve them, and a comparison with a classical solver is also possible.
+In particular, we address the train rescheduling problem in complex railway networks with mixed infrastructure including single, double, and multiple-track railway lines with given planned train paths.
 Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections.
-We apply a new hybrid rescheduling algorithm combining classical-quantum modelling and based on Quantum Annealing (QA).
-This work extends on a particular linear modeling strategy, partly explored on a toy model in [21].
-We develop an Integer Linear Programming (ILP) model, which is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
-The D-Wave approach is analysed in detail, with respect to its applicability and performance in a practically relevant situation.
-The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready for tackling challenging railway rescheduling problems.
-During the preparation of the present work, another contribution has appeared in the literature [56], which is similar to our in that it sets up a QUBO model for a somewhat similar problem, and solves it by a simulation of (quantum) annealing.
-This contribution, however, considers a different railway infrastructure: high-speed trains, and considers timetable optimization, for which it employs a model different from the one described here.
-The main contributions of the present work are the following:
-  • A railway traffic dispatching model is introduced for rescheduling trains due to disturbances and disruptions taking into account given timetable, rolling stock connections and a network with single or multi-track segments.
-  • QA-based hybrid heuristics are, for the first time, applied to real-life problems of railway rescheduling optimization on urban railway networks, these heuristics have some but still meaningfully input of real QA.
-  • The performance of the hybrid (quantum classical) D-Wave solver is demonstrated on a real-life network in Poland and diverse disturbance scenarios of different sizes and types.
-  • The D-Wave hybrid solver provides good quality solutions in short time, also for the complex instances in limited computational time.
-  • For certain instances the hybrid (quantum classical) D-Wave solver outperforms classical solvers in terms of computational time, yielding still feasible solutions.
+We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA).
+This work extends on a particular linear modeling strategy, partly explored on a toy model by Domino et al.
+(2022b).
+Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
+We recognize several important scientific gaps.
+First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances.
+Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices.
+Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling.
+In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks.
 The paper is organized as follows.
-In Section 2 we review state-of-the-art literature and identify scientific gaps we intend to fill, in Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
-2 Literature review
-In this section, we present time-indexed models known from the literature, comparing them briefly with our modelling approach.
-For a general review of railway timetabling and rescheduling we refer to the reviews by [38] and [7].
-Second, we review recent QA applications in optimization in general, and also in the rail/transport domain.
-Finally, we summarize the existing scientific gaps.
-2.1 Classical railway dispatching/rescheduling
-Time-indexed modelling of railway scheduling and rescheduling is quite common in literature for routing [39, 8] and scheduling [9, 29, 43, 13] .
-[8] focus on determining train’s path for a complex central railway station area (called the blocking - stairway), used a discrete-time model resulting in a $0-1$ program.
-Their model is successfully demonstrated on an operational day at the central railway station area Berne, Switzerland.
-Similarly, [39] consider train movements in a single major (and thus complex) railway junction, including a freight yard and a few minor stations.
-They build a mixed integer programming model that can also be solved close to optimal in case of practically relevant instances.
-They address the train movements on the level of detail of acceleration and deceleration strategies, hence adopting a significantly finer discrete time scale than our one-minute resolution.
-When compared to our problem, these two models addressed a high traffic in a complex station area.
-Instead, we are concerned more with a bigger area, multiple stations, and mixed track, but a lower train density and complexity.
-Our train routes fixed within the stations, and we also model shunting movements that is not a subject of the cited reference.
-[9] proposed a graph theoretic formulation for the problem using a directed multigraph in which nodes correspond to departures/arrivals at a certain station at a given time instant.
-This formulation is used to derive an integer linear programming model that is solved using a Lagrangian relaxation.
-[13] introduced a new pure 0-1 programming formulation, and call it Tick Formulation, to model the Deadlock Detection (DD) problem.
-[43] focused “on the train dispatching problem on an N-track network, with the main challenge of how to formulate specific retiming, reordering, retracking and rerouting options in combination”.
-Their model was demonstrated on generic networks of different structure than what we address.
-In our situation we are limited to retiming and reordering of trains.
-Structurally, their model shows similarities to ours, including the role of binary variables and also the introduction of "cells" that generalize the objects the occupancy of which can be controlled: they can be one or more blocks, stations, etc.
-In our model, the decision variables will be explicitly linked to a subset of stations.
-2.2 Quantum Annealing and its applications
-Quantum Annealing (QA) [15] is an optimization method analogous to Simulated Annealing (SA).
-Both can be used for the unconstrained minimization of a quadratic function of ±$\pm 1$-valued variables, which represents the energy configuration for a set of spins of the Ising model [30, 4]; an important and celebrated problem in physics.
-The Ising problem can be equivalently rewritten in a form of a Quadratic Unconstrained Binary Optimization (QUBO) problem with binary ($0-1$) variables.
-The Max-Cut problem [5] is also equivalent to the Ising model or QUBO.
-The problem is NP-hard, has tremendous literature in classical optimization, and there is significant ongoing progress in the development of algorithms to solve it [22, 28].
-An extensive list of Ising problems and their formulations are presented e.g.
-in the work of [37].
-Hardware quantum annealers, like the D-Wave machine, implement a quantum version of the Ising model, assigning a quantum bit, i.e.
-a two-level physical quantum system to each bit.
-The system employs real physical two-level quantum systems with a tunable energy operator (Hamiltonian), the energy being the objective function of the optimization problem.
-An adiabatic evolution of the quantum system is implemented: the interactions between the spins are slowly changed from an initial Hamiltonian with a simple minimal-energy state to the Hamiltonian corresponding to the objective according to the rules of quantum mechanics.
-According to the adiabatic theorem of quantum mechanics [3], under certain conditions the physical system remains in its lowest energy state during the evolution, and an optimal, minimum-energy solution can be possibly read out at the end.
-An ideal adiabatic quantum computer would be a device with a large number of perfect quantum bits completely decoupled from their environment, very close to zero temperature.
-Any pair of qubits could be coupled, and even, more general couplings, e.g.
-involving $3$ or more spins could take place.
-Such a system would exhibit perfect quantum coherence and entanglement.
-Even in such an ideal system the evolution time to reach at a minimum energy state with high enough probability is inverse proportional to the gap between the minimal energy configuration and the one closest to it.
-Certainly the gap is not known in advance, hence, the annealing time has to be determined by experimenting in practice.
-A more severe issue is that the gap can be very small , hence, it is not possible to solve all possible hard problems on such a setup.
-In spite of that, there are problems in which this approach can be efficient.
-[…]
-Operationally, from the optimizer’s point of view quantum annealers can be viewed as probabilistic heuristic solvers returning a statistical sample [20] of configurations which are supposed to be optimal or close to optimal.
-It must be stressed that the quantum annealers are not algorithms running on digital computers; they are analog devices implemented physically.
-This latter implies that the coefficients of the problem are encoded with a limited accuracy, and the algorithmic properties of the particular optimization problem such as its complexity class will not determine the solver’s behaviour.
-To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently.
-Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ [11], including the one used in the present work, belong to this family.
+In Section 2 we briefly review state-of-the-art of quantum annealing and its applications.
+In Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
+2 Quantum annealing
+Quantum computing devices became available for practical computational purposes in the last few years.
+Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics (Ising, 1925; Bian et al., 2010).
+The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
+Hence, a QUBO objective can be expressed as the energy of the Ising system: the configuration of the pairwise couplings of the spins and the applied local fields, which are tunable in quantum hardware.
+According to the adiabatic theorem of quantum mechanics (Avron and Elgart, 1999), tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out.
+Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems.
+Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  (Das and Chakrabarti, 2008).
+In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs.
+They output a number of configurations typically close to, and optionally including the optimal one.
+The class of QUBO problems is well studied and have a number of applications; we refer the reader to a recent monograph edited by Punnen (2022).
+Linear unconstrained binary problems can also be considered as a special case of QUBOs, while linear and quadratic constraints can be taken into account with penalties (Gusmeroli et al., 2022).
+Using appropriate binary encoding of the variables, integer programming problems can also be treated.
+Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
+The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled.
+The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system.
+This embedding is a hard problem itself (Zbinden et al., 2020).
+It often requires to couple multiple physical qubits to represent a logical bit of the problem.
+As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem are encoded with a limited accuracy.
+To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum-classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently.
+Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a), including the one used in the present work, belong to this family.
 Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
 Currently, the applicability of QA technologies is being explored.
-Benchmark problems are solved [55, 42], comparisons with digital computers are made [32], etc.
-The application of quantum annealers in transportation optimization is a new area with only a few contributions so far.
-There is an apparent interest in this research direction in the aviation domain [47, 51, 48], which relates also to the already mentioned benchmarking of annealers as done by[55].
-Shipment rerouting was also addressed in the context of QA [58].
-There are a few examples of hybrid solvers’ applications, too.
-These include traffic flow optimization [44], multi-car paint shop optimization [57], tail assignment problem [40], and vehicle optimization [24].
-In railway context, the first applications of QA can be found in train dispatching/rescheduling [19, 21] and rolling stock planning [27].
-In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [18].
-A followup paper [21] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems.
-The present work continues this line of research.
-2.3 Scientific gaps
-We recognize several important scientific gaps.
-First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances.
-Second, the QA models for the studied optmimzation problem that have been introduced so far [19, 21] had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices.
-Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling.
-In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks.
-To do so, we use a hybrid approach combining quantum and classical computing.
-For a fair comparison of current and future QA approaches with state-of-the-art classical approaches, we elaborate the railway model readily suitable for both classical and hybrid (quantum classical) solvers in a comparable manner.
-Quantum computing is believed to develop rapidly in near future.
-Hence, demonstration quantum readiness for railway problems right now is step forward towards the demonstration of quantum advantage in future.
-On future quantum devices the comparison with classical solvers is expected to be in strong favour for the quantum ones.
-In this way we also contribute in developing a novel general set of railway rescheduling models that has a potential to scale well with the size of the problem, aiming to overcome the curse of dimensionality.
+In the railway context, the first applications of QA can be found in train dispatching/rescheduling (Domino et al., 2022b) and rolling stock planning (Bickert et al., 2021).
+In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us (Domino et al., 2023).
+A follow-up paper (Domino et al., 2022b) laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems.
+The present work continues this line of research, which is gaining increasing attention, e.g.
+the recent work of Xu et al.
+(2023).
 3 Problem description
@@
 We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations.
-Also, we consider an urban railway network with mixed tracks from single, double, up to quadruple tracks.
+Also, we consider an urban railway network with mixed tracks from single, double, to quadruple tracks.
 We model a railway network with edges and nodes of a graph.
-Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple track lines).
+Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple-track lines).
 Each edge is composed of one or more tracks.
-Each track consists of blocks sections defined by pairs of signals [39].
+Each track consists of block sections defined by pairs of signals (Lusby et al., 2013).
 Each block section can be occupied by at most one train at a time.
 A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains.
-We assume a 2-block signalling system, meaning that two free blocks are required between the consecutive trains.
-We also assume green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
+We assume a 2-block signaling system, meaning that two free blocks are required between the consecutive trains.
+We also assume a green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
 Each node is composed of station tracks (blocks) and interlocking areas.
 One station track can be occupied by one train at a time.
-Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered in order to to guarantee that only one train from the pair can occupy the area at the time.
+Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered to guarantee that only one train from the pair can occupy the area at the time.
 A train’s route is the sequence of blocks the train passes during its journey.
 A train path is a sequence of arrival and departure times of a particular train assigned to a train route.
-We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time or passing time, resulting in integer variables.
+We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time, or passing time, resulting in integer variables.
 This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning the inherently binary variables.
-The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to precedence of trains, encoded later on precedence binary variables.
+The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to the precedence of trains, encoded later on precedence binary variables.
 In the network, two trains can follow each other, i.e.
@@
 change the order, when going the same direction.
-Single tracks are designed for bidirectional traffic, double-tracks for unidirectional, one for each direction, and multi-track lines can be combination of unidirectional and bidirectional.
+Single-track lines are designed for bidirectional traffic, double track lines for unidirectional, one for each direction, and multi-track lines can be a combination of unidirectional and bidirectional.
 On single-track lines, M-Ps and M-Os are only possible at stations.
-To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single track line heading in opposite directions.
+To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single-track line heading in opposite directions.
 Trains in the same direction preserve their order between stations, and keep the minimal headway time between each other.
 To prescribe these we determine set $\mathcal{J}^{2}_{\text{headway}}$: the set of all pairs of trains that can potentially violate the minimal headway condition.
-The train traffic is scheduled on the basis of the given timetable.
+The train traffic is scheduled based on the given timetable.
 The timetable contains all the train paths.
@@
 However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc.
-Following [10], a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously.
+Following Corman et al.
+(2012), a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously.
 We assume that the conflict may occur either on the railway line or at the station.
-Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single track line by two trains heading in opposite directions at the same time.
-Conflicts at stations include claiming a station track by two trains at the same time, claiming a station switch (in the interlocking area) by two trains at the same time.
-The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing and retiming for trains claiming for the same resource.
+Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single-track line by two trains heading in opposite directions at the same time.
+Conflicts at stations include claiming a station track by two trains at the same time, or claiming a station switch (in the interlocking area) by two trains at the same time.
+The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource.
 Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable.
 These changes typically result in delays of additional trains that were not directly involved in the conflict otherwise, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable.
-It is common in the literature to minimize a function of these secondary delays (see e.g.
-the work of [14]); we will do this in the present contriubution, too).
-An important element of our approach is that from amongst all stations we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only.
-Hence, as decision stations we select those stations, where routes of trains intersect, where trains start or terminate, or where selected part network is bounded.
-The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths has to be made with respect to these stations.
-(If we change some of the trains’ routes in a re-routing process, new model with new parameters if developed, e.g.
+We will minimize a function of these secondary delays.
+An important element of our approach is that from among all stations, we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only.
+Hence, as decision stations, we select those stations, where routes of trains intersect, where trains start or terminate, or where the selected part of the network is bounded.
+The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths have to be made with respect to these stations.
+(If we change some of the trains’ routes in a re-routing process, the new model with new parameters is developed, e.g.
 additional decision stations may appear.) In what follows, by a station we always mean a decision station.
@@
 The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa.
-The depot is treated as the station, and consider it as a black box, without detailed layout.
-We treat shunting movements as service train from depot to the starting station of the service train, or from the terminating station to the depot.
+The depot is treated as the station, and consider it as a black box, without a detailed layout.
+We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot.
 The rolling stock circulation condition is applied to ensure the precedence between the service train and the actual train.
@@
 4 Methods and Model
-In the following we describe our model in detail. Section 4.1 defines sets, parameters and decision variables. Section 4.2 describes our Integer Linear Programming (ILP) formulation.
+In the following, we describe our model in detail. Section 4.1 defines sets, parameters, and decision variables. Section 4.2 describes our Integer Linear Programming (ILP) formulation.
 4.1 Sets, parameters and decisions
-To formulate our decision variables, constraints and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan.
+To formulate our decision variables, constraints, and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan.
 Also, we introduce parameters calculated from the same input.
-4.1.1 Sets
+Sets
 Let us denote by $\hat{S}$ the set of all stations, and by $S\subset\hat{S}$ the decision stations we have determined in advance, as already described.
@@
 This set is deduced from the rolling stock circulation plan and the timetable.
-Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains which are close enough to each other in time so that precedence variables have to be defined for them.
-More details will be given at the description of the parameter $d_{\max}$ which this set depends on.
+Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains that are close enough to each other in time so that precedence variables have to be defined for them.
+More details will be given in the description of the parameter $d_{\max}$ which this set depends on.
 Set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that can potentially violate the minimal headway time between two subsequent trains on a line segment.
@@
 The order within the pairs is determined by $j$.
-All these sets can be enumerated on the basis of the input data in a straightforward manner.
-4.1.2 Parameters
+All these sets can be enumerated based on the input data in a straightforward manner.
+Parameters
 The parameters that appear in our model are the following:
@@
   • $\sigma(j,s)$ is the scheduled departure time of $s$ from $j$, given by the original timetable.
-  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance and/or disruption in the network happens.
+  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance in the network happens.
 It is the maximum of $\sigma(j,s)$ and the technically feasible earliest departure time, and may be more constraining than the first depending on the initial delays.
-The latter is calculated assuming that the train follows its planned route at minimum running time, and that there are no other trains present on the network.
+The latter is calculated assuming that the train follows its planned route at minimum running time and that there are no other trains present on the network.
   • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$.
-This parameter sets an upper limit for the possible secondary delays that can arise on the particular part of the network.
+This parameter sets an upper limit for the possible secondary delays that can arise on a particular part of the network.
 Setting such a bound is common in the literature, see e.g.
-the work of [14].
+the work of D’Ariano et al.
+(2007).
 We set the same value of this parameter for all our testing instances.
@@
 $d_{\max}=40$ excludes the possibility of a one-hour delay due to waiting for another delayed train).
-Meanwhile it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
+Meanwhile, it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
 The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $J^{2\,(\text{close})}$.
-A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed/disrupted timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$.
+A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$.
 The number of constraints will be proportional to the average number of trains that can meet another train at a station; this will be linear in $d_{\max}$.
-4.1.3 Decision variables
+Decision variables
 Our decision variables are the following. First, we use departure time variables:
@@
 The last set of precedence variables describes the precedence of trains at some resource located between stations $s$ and $s^{\prime}$ (e.g.
-single track used by $j$ and $j^{\prime}$ heading in opposite direction).
+a single-track segment used by $j$ and $j^{\prime}$ heading in opposite direction).
 The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$.
@@
 Given the index sets, variables, and the parameters and decision variables of the model, now we formulate the optimization objective and the constraints.
-4.2.1 Objective function
+Objective function
 Our goal is to minimize the secondary delays that occur in the analyzed part of the railway network.
@@
 The constant multiplier $1/d_{\text{max}}$ is optional; we use it for better comparability of different instances.
-4.2.2 Constraints
+Constraints
 The constraints of the model are the following.
@@
 where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$.
-(Constraints of this structure are often termed as big-M constraints in this context.) In our implementation we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
+(Constraints of this structure are often termed as big-M constraints in this context.) In our implementation, we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
 $C=-\min(t^{(\text{out})}(j^{\prime},s))+\max(t^{(\text{out})}(j,s))+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})$ (5)
@@
 Station tracks can be occupied by at most one train at a time:
-$\small{\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
+${\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
 where $C$ is chosen similarly to that in Eq.
@@
 Rolling stock circulation constraints
-These are introduced in order to bind the train with the shunting movement called also service train.
+These are introduced to bind the train with the shunting movement called also service train.
 If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
@@
 As for the scaling of our model, the number of time variables i.e.
-#$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  [21]).
-In order to obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations.
+#$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  Domino et al.
+(2022b)).
+To obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations.
 We assume that in average each train visits $\alpha\in(0,1)$ fraction of the stations, we have then
@@
 as a tighter bound.
-Suppose, that most of the network is composed mainly of double track lines.
+Suppose, that most of the network is composed mainly of double-track lines.
 Then there will be typically a single precedence variable per train and station ($y^{(out)}(j,j^{\prime},s)$), and thus their number can be estimated by:
 #$\#y+\#z\approx\#y\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S},$ (18)
-where $N(d_{\max})$ is number of trains the train can meet at a station on average. Under such assumptions the total number of variables grows linearly with number of trains and stations:
+where $N(d_{\max})$ is the number of trains the train can meet at a station on average. Under such assumptions, the total number of variables grows linearly with number of trains and stations:
 #$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}\#\mathcal{S}.$ (19)
-If, on the other hand, the network has dominantly single track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
+If, on the other hand, the network has dominantly single-track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
 #$\#y=\#z\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S}.$ (20)
@@
 Based on Eq.
-(16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single track occupation constraints.
-For the interlocking area occupation constraints it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable.
+(16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single-track occupation constraints.
+For the interlocking area occupation constraints, it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable.
 For the running time and minimal dwell time we expect one constraint per train and station.
-Thus number of constraints can be estimated by
+This number of constraints can be estimated by
 #$\#constr.\approx 6\#y+2\#z+2\#\mathcal{J}\#\mathcal{S}$ (21)
-where the second term is tied to single track line conditions (one condition per $z$ variable).
-#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single track}.\end{cases}$ (22)
+where the second term is tied to single-track line conditions (one condition per $z$ variable).
+#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double-track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single-track}.\end{cases}$ (22)
 Because of the limitations in time imposed by $d_{\max}$ the model can be viewed as linear in number of trains and number of stations.
-Hence, as expected, the model is more complex for the network in which single track lines dominate.
-Although the analysed network is mostly composed of the double track lines, we will analyse certain use-cases with the fraction of single track lines increased.
-Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables and number of constraints in the number of trains.
+Hence, as expected, the model is more complex for the network in which single-track lines dominate.
+Although the analyzed network is mostly composed of the double-track lines, we will analyze certain use-cases with the fraction of single-track lines increased.
+Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables, and that of the number of constraints in the number of trains.
 5 Hybrid quantum-based approach
-To overcome the limitations described in in Section II of current hardware quantum annealers, we apply hybrid (quantum classical) solver.
-In particular, we use the ’Leap Hybrid Solver Service (HSS)’ [11]; a cloud-based proprietary solver which is developed by the market leader of QA hardware and is available as a service.
+To overcome the limitations described in in Section 2 of current hardware quantum annealers, we apply a hybrid (quantum-classical) solver.
+In particular, we use the ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a); a cloud-based proprietary solver that is developed by the market leader of QA hardware and is available as a service.
 It utilizes a hybrid approach that combines classical computational power with quantum processing.
-In particular, we have used the Constrained Quadratic Model (CQM) [12] Solver. The CQM hybrid solver works as follows.
+In particular, we have used the Constrained Quadratic Model (CQM) (D-Wave Quantum Inc., 2022b) Solver. The CQM hybrid solver works as follows.
 Input
-A constrained linear or quadratic model, and the value of the solver parameter t_min, which prescribes the minimum required run time (in seconds) solver is allowed to work on any problem.
+A constrained linear or quadratic model, and the value of the solver parameter t_min, which prescribes the minimum required run time (in seconds) the solver is allowed to work on any problem.
 Preprocessing
-Preprocessing includes subproblem identification and decomposition to smaller instances [52].
+Preprocessing includes subproblem identification and decomposition to smaller instances (Tran et al., 2016).
 Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
@@
 The hybrid solver implements a workflow that combines a portfolio of various classical heuristics (including tabu search, simulated annealing, etc.) running on powerful CPUs (Central Processing Units) and GPUs (Graphics Processing Units).
-In the course of the solution process the hardware quantum annealer is also invoked, in order to solve smaller QUBO problems that can potentially boost the classical heuristics.
-Integer problems are transformed to binary using state-of-the-art methods like the one described by [25][34], generally applicable for integer problems.
+In the course of the solution process the hardware quantum annealer is also invoked, to solve smaller QUBO problems that can potentially boost the classical heuristics.
+Integer problems are transformed to binary using state-of-the-art methods like the one described by Glover et al.
+(2019); Karimi and Ronagh (2019), generally applicable for integer problems.
 When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions.
@@
 In this way, even though the physical annealer supports problems with limited size (e.g.
-the Pegasus machine has $5500$ physical qubits [16], and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process.
+the Pegasus machine has $5500$ physical qubits (Dattani et al., 2019), and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process.
 The final solution is composed by the solver’s module.
 Output
-The solution of the integer or quadratic constrained problem.
-The workflow results in a solution that is not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver.
-Furthermore, the solver is probabilistic, hence each at each run the final solution may be different.
+The solution or solutions of quadratic constrained problem or the integer problem.
+The workflow results in a solution or solutions that are not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver.
+Furthermore, the solver is probabilistic, hence at each run the final output can be different.
+The CQM (integer) solver returns a series of solutions, which can be an additional asset.
+Nevertheless, for sake of presentation, our approach is to analyze only the best CQM solution given each run of the hybrid algorithm (realization).
 While D-Wave’s proprietary solvers act as black boxes, i.e.
-they hide the exact details of the described process process from the user, the output is supplemented with timing parameters, including run time: the total elapsed time including system overhead, and Quantum Processing Unit (QPU) access time which is is the time spent accessing the actual quantum hardware.
+they hide the exact details of the described process from the user, the output is supplemented with timing parameters, including run time: the total elapsed time including system overhead, and Quantum Processing Unit (QPU) access time which is the time spent accessing the actual quantum hardware.
 These parameters enable a comparison with other solvers.
 As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min.
-This is a time limit for heuristics running in parallel[12]: unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution.
-We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting to the model performance and solution quality for our problem.
+This is a time limit for heuristics running in parallel(D-Wave Quantum Inc., 2022b): unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution.
+We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting on the model performance and solution quality for our problem.
 The CQM Solver can solve problems encoded in the form of Eq.
@@
 Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints.
-Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum classical) solvers.
+Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum-classical) solvers.
 As a classical solver, IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) was used.
 The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory.
-As hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.
+As a hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.
 6 Computational results
-In this section, we demonstrate the performance of the proposed hybrid (quantum classical) approach on a network at Polish railways - central part of the Upper Silesia Metropolis.
-In particular, we compare the performance of DWave’s hybrid quantum-classical solver (CQM) with CPLEX.
-The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances and disruptions including various initial delays and track closures.
-As the objective is tied to secondary delays, we can roughly apply it to asses the degree of difficulty of the problem.
+In this section, we demonstrate the performance of the proposed hybrid (quantum-classical) approach on a network at Polish railways - the central part of the Upper Silesia Metropolis.
+In particular, we compare the performance of D-Wave’s hybrid quantum-classical solver (CQM) with CPLEX.
+The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances including various initial delays and track closures.
+As the objective is tied to secondary delays, we can roughly apply it to assess the degree of difficulty of the problem.
 Section 6.1 describes the considered network and traffic characteristics.
-Section 6.2 presents results concerning generic examples on the core line of the network, comparing double track line scenarios with single track line scenarios and closure scenarios.
+Section 6.2 presents results concerning generic examples on the core line of the network, comparing double-track-line scenarios with single-track-line scenarios and closure scenarios.
 Section 6.3 presents analogous results for larger network with real-life timetable.
-Here the case studies concerns various scenarios with different degree of difficulty.
+Here the case studies concern various scenarios with different degrees of difficulty.
 6.1 Railway network
 We use the Open Railway Map [^1] https://www.openrailwaymap.org/, visited 2022-11-25 representation of the selected part of the Polish railway network located in the central part of the Metropolis GZM (Poland), presented in Fig 2.
-The selected part of the network combines heterogeneous network layout including single-, double-, triple-, and quadruple-track lines.
+The selected part of the network combines a heterogeneous network layout including single-, double-, triple-, and quadruple-track lines.
 It comprises 25 nodes including $11$ stations and $3$ junctions, $146$ blocks, and 2 depots.
@@
 Figure 2: Selected part of Polish railway network located in the central part of the Metropolis and examined in our research.
-The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analysed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O).
+The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analyzed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single-track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O).
 As described before, the decision stations are the only ones that appear explicitly in our model.
@@
 (2) we use $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable).
-As for the solver parameter t_min, after initial experiments we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality.
-Sensitivity analysis regarding this parameter is performed in Appendix A for selected instances.
+As for the solver parameter t_min, after initial experiments, we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality.
+A sensitivity analysis regarding this parameter will be performed in Section 6.3.
 6.2 Synthetic experiments on the selected railway line
 As the first set of experiments we address, synthetic instances on a part of the network - railway line KO-GLC, see Fig 2.
-The goal in these scenarios is to compare the CPLEX and CQM solvers for: the double track line ($line1$), the double track line with closures ($line2$), and the single track line ($line3$).
+The goal in these scenarios is to compare the CPLEX and CQM solvers for: the double-track line ($line1$), the double-track line with closures ($line2$), and the single-track line ($line3$).
 Shunting movements and rolling stock circulation are not considered in this set of experiments.
@@
   1. 1.
-    $Line1$, the double track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
+    $Line1$, the double-track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
   2. 2.
@@
   3. 3.
-    $Line3$, the whole line is considered as a single track, with a timetable of $3$ hours traffic and $21$ trains.
+    $Line3$, the whole line is considered as a single-track, with a timetable of $3$ hours traffic and $21$ trains.
 For $line1$ and $line3$, the original timetable is feasible, while for $line2$ it is not feasible due to the closure.
-For each of the lines we compute $12$ instances, each with different initial delays of trains subsets.
+For each of the lines, we compute $12$ instances, each with different initial delays of train subsets.
 Instance 00 yields no initial delays.
@@
 (17) - (19), (22) allows us to quickly estimate the sizes of the problem.
-This sizes are compared with computed (comp.) actual number of variables and constraints (mean over instances) in Tab.
+These sizes are compared with the computed (comp.) actual number of variables and constraints (mean over instances) in Tab.
 1.
-Here, $line2$ is the most complex in terms of constraints, what is reflected later by computational difficulties of some of its instances.
-Furthermore, $line2$ is partially double track and partially single track, hence the actual number of constraints is bounded from above by the estimate for the single track line and from below by the estimate for the double track line.
+Here, $line2$ is the most complex in terms of constraints, which is reflected later by the computational difficulties of some of its instances.
+Furthermore, $line2$ is partially double-track and partially single-track, hence the actual number of constraints is bounded from above by the estimate for the single-track line and from below by the estimate for the double-track line.
 $Line1$ includes double-, whereas $line3$ single-track segments only.
@@
 (17) - (19), (22) (doub.
-- double track, sing.
-- single track, comp.
+- double-track, sing.
+- single-track, comp.
 mean - computed actual number of variables or constraints, mean over instances).
-The computational results are presented in Figs.
-3,  4, and 5 for $line1$, $line2$, and $line3$, respectively.
+The computational results are presented in Fig.
+3 a) b) c) for $line1$, $line2$, and $line3$, respectively.
 Each figure shows objective $\times$ $d_{\text{max}}$ (top part), computation time (comp.
-time, middle part) and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s.
+time, middle part), and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s.
 CPLEX results are also displayed.
 For scenario $line1$ (Fig.
-3), CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$).
+3 a) ), the CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$).
 Similarly, for $line3$ (Fig.
-5), only small deviations from the optimum can be observed for a number of instances.
+3 c) ), only small deviations from the optimum can be observed for a number of instances.
 All computational times for $line1$ and $line3$ were significantly below $0.4$min, although CPLEX performed faster.
@@
 For $line2$ (Fig.
-4), all CQM solutions were suboptimal, although feasible.
+3 b) ), all CQM solutions were suboptimal, although feasible.
 More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than CPLEX, in several cases up to 5×5$\times$5 × faster.
 This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5).
-This observation is approved by more detailed discussion on t_min parameter in Appendix A.
 In principle, $line1$ (Fig.
-3) and $line3$ (Fig.
-5) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster.
+3 a) ) and $line3$ (Fig.
+3 c) ) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster.
 We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems.
-Figure 3: Synthetic experiments $Line1$, comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM.
+Figure 3: Synthetic experiments $Line1$, $Line2$, $Line3$ comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM.
 All the displayed solutions are feasible.
 Total computational time (middle panel) and QPU times (lower panel) were provided by the D-Wave output.
-Results of both solvers were similar in terms of the objective, but the CPLEX computation was faster.
-Figure 4: Synthetic experiments $Line2$, comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM.
-Note that the CQM is fast, even when compared to CPLEX.
-The found solutions are suboptimal but can be useful.
-All the displayed solutions are feasible.
-$Line2$, namely the double track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective).
+Panel a) - $line1$, results of both solvers were similar in terms of the objective, but the CPLEX computation was faster.
+Panel b) - $line2$ namely the double-track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective).
 On this example, there are instances, where the current CQM D-Wave hybrid solver outperforms CPLEX in terms of computational time.
 Increasing the value of the t_min parameter in most cases improves the quality of the solution at the cost of computational time.
-Figure 5: Synthetic experiments $Line3$, comparison of the performance of classical solvers (CPLEX) with that of the hybrid CQM.
-All the displayed solutions are feasible.
+Panel c) - $line3$ all the displayed solutions are feasible.
 Results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
 6.3 Real-life experiments on the considered rail network
-In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on presented network.
-We will also focus on the track closure situation, turning a double track line segment into a single track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2.
+In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on the presented network.
+We will also focus on the track closure situation, turning a double-track line segment into a single-track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2.
 Our computations address various use cases of train delays outside the network and within the network as well as the partial track closure.
-We constructed $9$ networks with increased level of difficulty ranging from small delays to disruptions, i.e.
+We constructed $9$ networks with increased levels of difficulty ranging from small delays to disruptions, i.e.
 it is expected more trains are involved and/or disturbances are spread more over the network.
@@
     1. (a)
-      Networks $4$ and $5$ concerns rerouting trains from double track line KTC-CB to the single track line with higher passing times, see Fig.
+      Networks $4$ and $5$ concerns rerouting trains from double-track line KTC-CB to the single-track line with higher passing times, see Fig.
 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
@@
       Network $6$ concerns multiple closures, i.e.
-change of multiple line KZ-KO(STM) and double track line KO-KL to single track lines.
+change of multiple line KZ-KO(STM) and double-track line KO-KL to single-track lines.
 (See Fig.
@@
 network #$\#S$ place trains involved description
-A. Disturbtions from outside the network
+A. Disturbances from outside the network
 1 10 KO-Ty IC 14006 KS 94113 Delayed IC (higher priority) is in conflict with on-time KS
 2 10 arr. Ty KS 94766 KS 40518 IC 41004 KS 44862 IC 4120 Delayed 5 trains on the arrival to Ty, random delays of several minutes
-B. Disturbtions originated within the network
+B. Disturbances originated within the network
 3 10 whole network 14 trains KS + IC Random delays of 14 trains delays up to 35 minutes
-C. Disturbtions within the network and closures
-4 11 KTC-CB 10 trains KS + IC Double track line KTC-CB closed trains rerouted to single track line
-5 11 whole network most trains Double track line KTC-CB closed as in 4 + delays as in 3
-6 10 whole network all trains Multiple track KZ - KO(STM) and double track KO - KL changed into single track + delays as in 3
-7 11 whole network all trains Multiple track KZ - KO(STM) and double track KO - KL changed into single track + double track line KTC-CB closed as in 4 + delays as in 3
+C. Disturbances within the network and closures
+4 11 KTC-CB 10 trains KS + IC double-track line KTC-CB closed trains rerouted to single-track line
+5 11 whole network most trains double-track line KTC-CB closed as in 4 + delays as in 3
+6 10 whole network all trains multiple-track KZ - KO(STM) and double-track KO - KL changed into single-track + delays as in 3
+7 11 whole network all trains multiple-track KZ - KO(STM) and double-track KO - KL changed into single-track + double-track line KTC-CB closed as in 4 + delays as in 3
 8 11 whole network all trains Closures as in 7 + random delays of 13 trains up to 30 minutes
@@
 Networks of particular dispatching problems.
-In networks 00 - $5$ most of analysed network is double tracked, from network $6$ upward the number of single track lines in the network increases at the expense of the number of double and multiple track lines.
+In networks 00 - $5$ most of analyzed network is double-tracked, from network $6$ upward the number of single-track lines in the network increases at the expense of the number of double and multiple-track lines.
 For each network we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains.
@@
 CQM hyb. t_min = $5$ s
-mean value over $5$ realis.
+mean value over $5$ realiz.
 comparison
@@
 We have selected t_min = $5$s as it provides the best balance between computational time and the quality of solution.
-For detailed analysis see Appendix A.
-This appendix shows a detailed analysis of the impact of t_min on computation time and solution quality for two specific real-life lines $7$ and $9$.
 The last columns present a percentage comparison (CQM - CPLEX ) / CPLEX, hence, the positive values represent CQM advantage, whereas negative values show CPLEX advantage.
@@
 This also facilitates the analysis of the statistics of the output and the differences between particular realizations.
-The level of difficulty increases with respect to the number of variables and constraints.
+The level of difficulty increases with the number of variables and constraints.
 Observe that CQM always returns a feasible solution.
@@
 As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$).
-Meanwhile those of CQM remain almost constant around $5.1s$.
+Meanwhile, those of CQM remain almost constant around $5.1s$.
 QPU times remain not greater than $0.032s$.
@@
 For example, in network $9$, the computed CQM solution is around 30% worse than the optimal one.
-In conclusion, we expect benefits from the application of the hybrid solver on medium scale railway networks with multiple closures, as in networks 7-9.
-This is in accordance with our observation on synthetic data in Fig.
-4, where the benefits appeared for more complex instances, too.
-To evaluate the extent of difficulty of the problems analysed in Tab.
+In conclusion, we expect benefits from the application of the hybrid solver on medium-scale railway networks with multiple closures, as in networks 7-9.
+This finding follows our observation on synthetic data in Fig.
+3 b), where the benefits appeared for more complex instances, too.
+To evaluate the extent of difficulty of the problems analyzed in Tab.
 3, Tab.
 4 presents mean (over stations and realizations) delays, resulting from CQM solutions, on selected stations (i.e.
-stations in central part of the network in Fig.
+stations in the central part of the network in Fig.
 2).
@@
 Fig.
-6 shows more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (top panel) vs CQM (bottom panel).
-For network $4$ (left panel of Fig.
-6), bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller.
+4 shows a more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (left, red) vs CQM (right, blue).
+For network $4$, bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller.
 This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Tab.
 2), and thus the spread is limited to these 3 stations.
-For network $7$ (right panel of Fig.
-6), delays are more spread through the network due to disruption and initial disturbances.
-To conclude, we can expect advantage from a hybrid (quantum classical) approach particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network.
+For network $7$, delays are more spread through the network due to disruption and initial disturbances.
+To conclude, we can expect an advantage from a hybrid (quantum-classical) approach, particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network.
+To facilitate further evaluation of the solutions, train diagrams are presented in Appendix A.
 Delay [min]
@@
 2.
-Non-zero values for network 00 comes from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes delay but make it up and leaves the network on time - these are typical issues for the railway traffic on analysed network.
+Non-zero values for network 00 come from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes, delay but makes it up and leaves the network on time - these are typical issues for the railway traffic on analyzed network.
 Figure 4: Real-life experiments.
-Statistics of CPLEX (top panel) and CQM (bottom panel) solutions (Tab.
+Statistics of CPLEX (left, read) and CQM (right, blue) solutions (Tab.
 3) for selected networks in Tab.
@@
 Vertical lines are ranges, while horizontal lines are $0.25$ and $0.75$ percentiles.
-In order to further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Figs.
-7,  8, and 9, respectively.
+The right choice of the t_min parameter can improve results meaningfully.
+Let us analyze in more detail the role of the t_min solver parameter for selected problems.
+In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig.
+5 a) and b).
+The objective value is roughly log-linear in computational time.
+For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially.
+Overall, the improvement in the objective is overwhelmed by the increase of computational time.
+This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling.
+Such behavior is plausible in the case of QA methods (Soriani et al., 2022; Albash et al., 2017; Domino et al., 2022a).
+Figure 5: Real-life experiments.
+The sweep over t_min parameter for network $7$ a) and network $9$ b), see Tab.
+2.
+All solutions were feasible.
+For higher t_min we expect smaller objective values but higher total computational time.
+For each parameter value, there were $5$ realizations of the experiment were performed.
+Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
+7 Conclusions
+We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks.
+In particular, we have encoded rescheduling problems in a linear integer program and solved these using a state-of-the-art solver (CPLEX), as well as the QA based CQM hybrid (quantum-classical) solver.
+The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems.
+While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better.
+This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
+While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models.
+In each computation, the CQM hybrid solver always has taken some advantage of the QPU.
+Hence, such a little help from the QPU must have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
+Several future research directions can be determined.
+The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices (DiVincenzo, 2000) with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular).
+The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of quantum device to our advantage e.g.
+by modeling some stochastic behavior on railways.
+The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation (Goto, 2019).
+With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.
+Data availability
+The code and the data used for generating the numerical results can be found in https://github.com/iitis/railways_dispatching_silesia under DOI 10.5281/zenodo.10657323
+Acknowledgments
+This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No.
+2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no.
+BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK).
+This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD).
+For the purpose of Open Access, the authors has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
+We would like to thank Sebastian Deffner as well as the whole scientific community of the Department of Physics, University of Maryland, Baltimore County (UMBC) for valuable discussion; and to Bartłomiej Gardas, and Zbigniew Puchała for valuable motivation; and Katarzyna Gawlak, Akash Kundu, and Özlem Salehi for data validation and assistance with coding.
+We acknowledge the cooperation with Koleje Ślaskie sp.
+z o.o.
+(eng.
+Silesian Railways).
+Appendix A. Train diagrams of CPLEX and CQM
+To further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Fig.
+6 (left CPLEX, middle and right two realizations of CQM).
 We have chosen the line segment between GLC and KZ, offering a good demonstration.
-(We have introduced a small artificial ’distance’ between KO and KO(STM) wich are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution.
+(We have introduced a small artificial ’distance’ between KO and KO(STM) which are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution.
 All the three proposed solutions have similar figures of merit, they all provide a valid option to be realized.
@@
   • in the CQM solution, second realization, 26103 passes first KTC-CB closure, then 5312 followed by 40150 go in opposite directions.
-This results in following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$.
-Although, the first solution is optimal, it is reasonable to supply dispatchers with the range feasible solutions.
-Hence, these diagrams proved that CQM hybrid solver can produce usable solutions for railway traffic management.
+This results in the following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$.
+Although the first solution is optimal, it is reasonable to supply dispatchers with the range of feasible solutions.
+Hence, these diagrams proved that the CQM hybrid solver can produce usable solutions for railway traffic management.
 An additional benefit of the probabilistic nature of the quantum-based method is that it produces different alternatives after each realization run without any modification.
@@
 The dispatcher’s choice can be influenced by other circumstances not covered by the present model.
-Figure 7: Real-life experiment. Time-distance diagram CPLEX solution of network $7$, see Tab. 2.
 Figure 6: Real-life experiment.
+Time-distance diagram solution of network $7$ CPLEX (left) and two realizations of CQM (middle and right), see Tab.
+2.
+Real-life experiment.
 Time-distance diagram CQM solution of network $7$ (see Tab.
 2) first realization.
-Closures (yielding single track line) are between CB and KTC as well as between KO(STM) and KZ.
-Figure 9: Real-life experiment. Time-distance diagram CQM solution of network $7$ (see Tab. 2) second realization.
-7 Conclusions
-We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks.
-In particular we have encoded rescheduling problems in a linear integer program, and solved these using a state-of-the-art solver (CPLEX), as well as on the QA based CQM hybrid (quantum classical) solver.
-While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better.
-This supports the expectation that future quantum devices will be efficient in solving large scale-problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that are beyond the scope of current exact models and heuristics.
-While the quantum-based solvers would possibly return suboptimal solutions, these are still often better than the solutions obtained manually or based on smaller scale models.
-More importantly, all generated solutions were feasible.
-In each computation, the CQM hybrid solver always has taken some advantage of the QPU.
-Hence, we believe that such a little help from the QPU must have boosted the classical heuristics in the hybrid solver.
-Recall that quantum annealer devices are subjects of many ongoing debates.
-Previous studies claimed that the quantum nature of their operation is limited [45], while recent works argue that their operation relies more on the thermalization with a cold environment in place of actual quantum adiabatic evolution [6].
-They are often criticized for the small size of problems they can address and various classical solvers do outperform them on certain problems at the present state of the art.
-In spite of all these, our computational results demonstrate the readiness of hybrid (quantum classical) solvers to handle real-life railway problems.
-We believe that future investigations shall focus on hybrid approaches both for railway applications as well as more generally.
-Finally, several future research directions can be determined.
-First concerns the ongoing worldwide works on the development of fault tolerant quantum computing devices [17] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular).
-Second concerns creation of custom open source hybrid solver that is dedicated for railway problems and will use noise of quantum device to our advantage e.g.
-by modelling some stochastic behaviour on railways.
-Third concerns evolution and application of quantum inspired techniques such as tensor networks technique, simulated bifurcation [26].
-With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.
-Data availability
-The code and the data used for generating the numerical results can be found in https://github.com/iitis/railways_dispatching_silesia under DOI 10.5281/zenodo.10657323
-Acknowledgements
-This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No.
-2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no.
-BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK).
-This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD).
-For the purpose of Open Access, the authors has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.
-We would like to thank to Sebastian Deffner as well as to the whole scientific community of the Department of Physics, University of Maryland, Baltimore County (UMBC) for valuable discussion; and to Bartłomiej Gardas, and Zbigniew Puchała for valuable motivation; and to Katarzyna Gawlak, Akash Kundu, and Özlem Salehi for data validation and assistance with coding.
-We acknowledge the cooperation with Koleje Ślaskie sp.
-z o.o.
-(eng.
-Silesian Railways).
-Compliance with ethical standards
-Conflict of interest
-The authors declare that they have no conflict of interest.
-Ethical approval
-This article does not contain any studies with human participants or animals performed by any of the authors.
-Appendix A. Comparative analyses for different values of the t_min parameter on selected solutions
-The right choice of the t_min parameter can improve results meaningfully.
-Hence, in this Appendix we analyse in more detail the role of the t_min solver parameter for selected problems in Section 6.3.
-In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig.
-10 and Fig.
-11.
-The objective value is roughly log-linear in computational time.
-For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially.
-Overall, the improvement in the objective is overwhelmed by the increase of the computational time.
-This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling.
-Such a behaviour is plausible in the case of QA methods [46, 1, 20].
-Figure 10: Real-life experiments.
-The sweep over t_min parameter for network $7$, see Tab.
-2.
-All solutions were feasible.
-For higher t_min we expect smaller objective values but higher total computational time.
-For each parameter value there were $5$ realizations of the experiment were performed.
-Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
-Figure 11: Real-life experiments. The sweep over t_min parameter for network $9$, see Tab. 2. All solutions were feasible, results were similar to those in Fig. 10.
+Closures (yielding single-track line) are between CB and KTC as well as between KO(STM) and KZ.

### v5

@@
 We address the applicability of hybrid quantum-classical heuristics for practical railway rescheduling management problems.
-We build an integer linear model for the given problem and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison.
-The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments and covers all the requirements posed by the operator of the network.
+We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison.
+The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments.
+All the requirements posed by the operator of the network are modeled.
 The computational results demonstrate the readiness for application and benefits of quantum-classical hybrid solvers in the realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation.
-At the same time, the solutions that were obtained were feasible.
-Moreover, though they are probabilistic (heuristics) they offer a valid alternative by returning a range of possible solutions the dispatcher can choose from.
-And, most importantly, they outperform classical solvers in some cases.
+In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
 Keywords
-railway rescheduling, conflict management, heterogeneous urban railway network, quantum annealing, hybrid quantum-classical heuristics
+[…]
 1 Introduction
-Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policy, leading to traffic challenges in passenger and cargo rail transport.
+Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policies, leading to traffic challenges in passenger and cargo rail transport.
 The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems.
-Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays (Ge et al., 2022).
-Examples of such disturbances include: late train departures/arrivals, extended dwell times.
-These can last from several minutes up to hours.
-A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too (Cacchiani et al., 2014).
-In the present work we focus on rescheduling of trains, which is the network operator’s perspective, even though we also consider (partial) track closures, which are are typically considered as disruptions.
-In our examples, however they will be managed by modifying existing train paths.
+Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [18].
+Examples of such disturbances include late train departures/arrivals, extended dwell times, etc.
+Their duration can be from several minutes up to hours.
+A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too [5].
+In the present work, we focus on train rescheduling, which is the network operator’s perspective.
+Even though we also consider (partial) track closures, which are typically considered as disruptions, in our examples they will be managed by modifying existing train paths.
 The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f.
-Törnquist (2007) and references therein).
+[37] and references therein).
 Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
-As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for efficient suppression of the consequences of disturbances.
-While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays (Lange and Werner, 2018), the maximal delay that cannot be avoided (D’Ariano et al., 2007), or fuel consumption measures  (Harrod, 2011).
-Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost real time.
-The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints (Szpigel, 1973; Mascis and Pacciarelli, 2002).
-It has an extensive literature, already summarized in a number of reviews, including those by Cacchiani et al.
-(2014), by Lamorgese et al.
-(2018), or by Philip and Swapnesh (2022).
-The maybe most frequently applied modeling strategy is based on alternative graphs (D’Ariano et al., 2007; Mascis and Pacciarelli, 2002) which can be formulated as as a mixed integer linear program, often referred to as the big-M model.
-As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and the combinations of these; see the work of Lamorgese et al.
-(2018) for a review of the model and the methods.
+As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for the efficient suppression of the consequences of disturbances.
+While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [27], the maximal delay that cannot be avoided [10], or fuel consumption measures  [22].
+Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost in real time.
+The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints [36, 29].
+It has an extensive literature, already summarized in a number of reviews, including those by [5], by [26], or by [31].
+The maybe most frequently applied modeling strategy is based on alternative graphs [10, 29], resulting in a disjunctive program.
+Such a model can also be converted into a mixed integer linear program referred to as the big-M model.
+As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and combinations of these; see the work of [26] for a review of the model and the methods.
 An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants.
-Even though this results in very large problems, it is applied for both timetabling and dispatching/rescheduling (Caimi et al., 2012; Lusby et al., 2013; Meng and Zhou, 2014).
-As this approach results in a $0-1$ program, which is promising when one considers the use of $0-1$ solvers such as quantum hardware.
-Indeed, job-shop problems addressed on pure quantum hardware (Venturelli et al., 2015), and even in certain hybrid approaches (Kurowski et al., 2020) use similar models, and our previous work (Domino et al., 2023) which was the first to apply a quantum solver to railway dispatching problems also followed this route.
+Even though the so-arising $0-1$ programs can be large, this method has been applied for both timetabling and dispatching/rescheduling [6, 28, 30].
+Meanwhile, $0-1$ programs are suitable as input for certain quantum hardware types.
+Indeed, job-shop problems addressed on hardware quantum annelaers [39], and even in certain hybrid approaches [25] use similar models, and our previous work [15] which was the first to apply a quantum solver to railway dispatching problems also followed this route.
 In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP).
@@
 Our consideration includes shunting movement of rolling stock between depots and stations followed by rolling stock connections.
+We do not include train cancellations in the model.
 We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA).
-This work extends on a particular linear modeling strategy, partly explored on a toy model by Domino et al.
-(2022b).
+This work extends on a particular linear modeling strategy, partly explored on a toy model by [17].
 Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
 We recognize several important scientific gaps.
-First, existing optimization models suffer from the curse of dimensionality, and inability to solve larger real-life instances.
+First, existing optimization models do not scale well: they can become huge and inefficient for larger real-life instances.
+Exploiting the characteristics of a given railway scenario, one can possibly construct simplified models whose size remains small enough to be addressed with a suitable exact or heuristic solver, yet they capture the relevant optimization objectives and constraints.
+For the particular railway scenario and solver studied in the present paper, no such model is known.
 Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices.
-Third, no hybrid QA-based models have been used for real-life railway, or even other schedule-based modes, like public transport and air traffic planning and/or rescheduling.
-In this paper we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of a urban railway networks.
+Third, no hybrid QA-based models have been used for real-life railway optimization so far, nor in other schedule-based modes, like public transport or air traffic planning/rescheduling.
+In this paper, we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of an urban railway network.
 The paper is organized as follows.
-In Section 2 we briefly review state-of-the-art of quantum annealing and its applications.
-In Section 3 we describe problem under investigation, in Section 4 we evaluate our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
+In Section 2 we briefly review the state of the art of quantum annealing and its applications.
+In Section 3 we describe the problem under investigation, in Section 4 we present our model, in Section 5 we discuss the hybrid solver we use, in Section 6 we present computational results.
+In Section 7 conclusions are drawn.
 2 Quantum annealing
 Quantum computing devices became available for practical computational purposes in the last few years.
-Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics (Ising, 1925; Bian et al., 2010).
+Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics [23, 3].
 The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
-Hence, a QUBO objective can be expressed as the energy of the Ising system: the configuration of the pairwise couplings of the spins and the applied local fields, which are tunable in quantum hardware.
-According to the adiabatic theorem of quantum mechanics (Avron and Elgart, 1999), tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out.
-Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems.
-Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  (Das and Chakrabarti, 2008).
-In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs.
-They output a number of configurations typically close to, and optionally including the optimal one.
-The class of QUBO problems is well studied and have a number of applications; we refer the reader to a recent monograph edited by Punnen (2022).
-Linear unconstrained binary problems can also be considered as a special case of QUBOs, while linear and quadratic constraints can be taken into account with penalties (Gusmeroli et al., 2022).
+The class of QUBO problems is well studied and has a number of applications; we refer the reader to a recent monograph edited by [32].
+Linear binary problems can also transformed to QUBOs; linear and quadratic constraints can be taken into account with penalties [21].
 Using appropriate binary encoding of the variables, integer programming problems can also be treated.
 Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
+Expressing a QUBO as the energy of an Ising system one gets a configuration of pairwise couplings of spins and local fields applied to individual spins, which can be realized in tunable quantum hardware.
+According to the adiabatic theorem of quantum mechanics [2], tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out.
+Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems.
+Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  [11].
+In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs.
+They output a number of configurations typically close to, and possibly including the optimal one.
 The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled.
 The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system.
-This embedding is a hard problem itself (Zbinden et al., 2020).
+This embedding is a hard problem itself [41].
 It often requires to couple multiple physical qubits to represent a logical bit of the problem.
-As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem are encoded with a limited accuracy.
-To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum-classical) solvers, orchestrating classical algorithms and using QA as a subroutine in order to address hard problem instances more efficiently.
-Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a), including the one used in the present work, belong to this family.
+As quantum annealers are not algorithms running on digital computers but analog devices implemented physically, the coefficients of the problem can only be set with limited accuracy.
+To overcome the problem of limited size and accuracy, quantum annealers of the present state of the art are often used in hybrid (quantum-classical) solvers, orchestrating classical algorithms and using QA as a subroutine to address hard problem instances more efficiently.
+Solvers available in D-Wave’s ’Leap Hybrid Solver Service (HSS)’ [8], including the one used in the present work, belong to this family.
 Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
 Currently, the applicability of QA technologies is being explored.
-In the railway context, the first applications of QA can be found in train dispatching/rescheduling (Domino et al., 2022b) and rolling stock planning (Bickert et al., 2021).
-In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us (Domino et al., 2023).
-A follow-up paper (Domino et al., 2022b) laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems.
+In the railway context, the first applications of QA can be found in train dispatching/rescheduling [17] and rolling stock planning [4].
+In particular, the first proof-of-concept demonstration of a (pure) quantum computing approach to railway rescheduling was presented by some of us [15].
+A follow-up paper by [17] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems.
 The present work continues this line of research, which is gaining increasing attention, e.g.
-the recent work of Xu et al.
-(2023).
+the recent work of [40].
 3 Problem description
@@
 We consider a railway rescheduling problem that includes train reordering, retiming, and shunting movements with rolling stock circulation at stations.
-Also, we consider an urban railway network with mixed tracks from single, double, to quadruple tracks.
+The considered traffic takes place in an urban railway network with mixed track systems: segments of single, double, to quadruple tracks.
 We model a railway network with edges and nodes of a graph.
-Figure 1 depicts an exemplary network layout composed of nodes (station or junction) and edges (single, double, multiple-track lines).
+Figure 1 depicts an example of a network layout composed of nodes (station or junction) and edges (single, double, multiple-track lines).
 Each edge is composed of one or more tracks.
-Each track consists of block sections defined by pairs of signals (Lusby et al., 2013).
+Each track consists of block sections defined by pairs of signals [28].
 Each block section can be occupied by at most one train at a time.
 A subsequent train is allowed to enter the block only after a minimal headway: minimal time span between the trains.
-We assume a 2-block signaling system, meaning that two free blocks are required between the consecutive trains.
-We also assume a green way policy (trains have the free way to move at the maximal allowed speed between stations), resulting in constant running times for each block.
+To place the model into a real-life railway context, we adopt a 2-block signaling system, meaning that two free blocks are required between the consecutive trains.
+The 2-block signaling system is common across the Polish railway network, and typical for conventional primary railway lines.
+The signaling system affects the model through the way of calculating parameter values (headway times), hence, the model can be applied to various safety systems.
+We assume green way policy, that is, trains have the free way to move at the maximal allowed speed between stations.
+This results in constant running times for each block.
 Each node is composed of station tracks (blocks) and interlocking areas.
-One station track can be occupied by one train at a time.
-Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered to guarantee that only one train from the pair can occupy the area at the time.
+The conflicts described by our model are the following.
+Each station track can be occupied by at most one train at a time.
+Routing dependencies between pairs of trains competing for the same resource in interlocking areas are considered to guarantee that only one train from the pair can occupy the area at a time.
 A train’s route is the sequence of blocks the train passes during its journey.
@@
 We assume a one-minute resolution for all time parameters such as timetable time, running and dwell time minimal headway time, or passing time, resulting in integer variables.
-This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning the inherently binary variables.
-The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to the precedence of trains, encoded later on precedence binary variables.
-In the network, two trains can follow each other, i.e.
-keep the given order, meet and pass (M-P) when going the opposite directions, and meet and overtake (M-O), i.e.
-change the order, when going the same direction.
-Single-track lines are designed for bidirectional traffic, double track lines for unidirectional, one for each direction, and multi-track lines can be a combination of unidirectional and bidirectional.
+This assumption of discretized time is needed to enable the use of quantum-based solvers, and, more importantly, introduces the possibility of pruning binary variables.
+The relaxation of integer constraints on time variables is not expected to improve computation time significantly, as the real difficulty is tied to the precedence of trains which is modeled by the binary precedence variables.
+In the network, two trains can follow each other keeping their given order.
+They can meet and overtake (M-O) when they go in the same direction, thereby changing their order.
+A pair of trains heading in the opposite direction can meet and pass (M-P).
 On single-track lines, M-Ps and M-Os are only possible at stations.
-To prevent M-P on single-track lines, we determine the set $\mathcal{J}^{2}_{\text{single}}$ as the set of all pairs of trains that can potentially meet on the single-track line heading in opposite directions.
-Trains in the same direction preserve their order between stations, and keep the minimal headway time between each other.
-To prescribe these we determine set $\mathcal{J}^{2}_{\text{headway}}$: the set of all pairs of trains that can potentially violate the minimal headway condition.
+This has to be enforced by constraints in our model.
+To maintain a reasonable number of constraints, the set of all pairs of trains that are involved in such constraints, that is, which can potentially meet on any single-track line segment in opposite directions, can be determined in advance from the model instance data.
+Similarly, trains in the same direction preserve their order between stations and keep the minimal headway time between each other.
+This does not hold for trains running in the same direction on multiple-track segments on different tracks.
+The set of all pairs of trains that can potentially violate the minimal headway condition can also be predetermined, thereby minimizing the number of the respective constraints.
 The train traffic is scheduled based on the given timetable.
@@
 However, conflicts may appear due to disturbances such as late departures and/or arrivals due to excessive passenger demand, malfunctioning rolling stock, etc.
-Following Corman et al.
-(2012), a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously.
-We assume that the conflict may occur either on the railway line or at the station.
-Possible conflicts that can occur on the line include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single-track line by two trains heading in opposite directions at the same time.
-Conflicts at stations include claiming a station track by two trains at the same time, or claiming a station switch (in the interlocking area) by two trains at the same time.
+The disturbances appear in our model via initial delays: the delays of trains at the beginning of their journey through the network.
+The initial delays have fixed values in each of our problem instances; the study of the statistics of delays because of random initial delays [35] is beyond the scope of the present study.
+Following [7], a conflict is an inadmissible situation when at least a pair of trains claim the same resource (e.g., block section, switch) simultaneously.
+Conflicts can occur either on railway lines or at the stations.
+Possible conflicts that can occur on the lines include the lack of the minimum headway between two subsequent trains heading on the same track in the same direction, or claiming a segment of a single-track line by two trains heading in opposite directions at the same time.
+Conflicts at stations include claiming a station track by two trains at the same time or claiming a station switch (in the interlocking area) by two trains at the same time.
 The conflicts have to be solved by modifying the original timetable, applying decisions on the train sequencing, and retiming for trains claiming the same resource.
 Such an intervention in the structure of train paths implies additional changes to maintain the feasibility of the modified timetable.
-These changes typically result in delays of additional trains that were not directly involved in the conflict otherwise, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable.
-We will minimize a function of these secondary delays.
-An important element of our approach is that from among all stations, we first determine a subset of the decision stations; we assume that the direct decisions implied by our model affect these stations only.
-Hence, as decision stations, we select those stations, where routes of trains intersect, where trains start or terminate, or where the selected part of the network is bounded.
-The motivation is that if the routes of trains are fixed, the decisions on modifying their train paths have to be made with respect to these stations.
-(If we change some of the trains’ routes in a re-routing process, the new model with new parameters is developed, e.g.
-additional decision stations may appear.) In what follows, by a station we always mean a decision station.
-This also means that non-decision stations appear only through the parameter values in the model, they do not appear as indices of decision variables or parameters.
+These changes typically result in additional delays of trains, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable.
+We will minimize the function of these secondary delays.
+First we determine the subset of decision stations: these are the nodes of the network (typically: stations) where the events that are directly affected by the decisions can take place.
+These events are: starting, termination, entering or leaving the modeled part of the network, meet and pass, meet and overtake, and the joining and splitting of train routes.
+The choice of decision stations depends on railway operation practices and actual traffic circumstances: stations where trains are allowed to meet and pass or meet and overtake, network junctions, and the boundaries of the selected part of the network have to be considered as decision stations.
+Meanwhile, there can be stations with an infrastructure supporting meet and pass or meet and overtake, which are not commonly used for this purpose under normal circumstances.
+These are normally not considered as decision stations.
+In the case of disruptions, however, some of them can be introduced as additional decision stations upon the network operator’s decision.
+From the model’s point of view, the decision stations are those to which decision variables are tied to.
+In what follows, we use the term "station" for decision stations only, unless otherwise stated.
+The model does not include decision variables tied to non-decision stations.
+For instance, the precedence of trains cannot be changed on non-decision stations.
+They do have their impact on parameter values in the model though.
 Headways, for instance, are calculated between decision stations, taking into account all the line blocks and station blocks of non-decision stations in between.
@@
 We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot.
-The rolling stock circulation condition is applied to ensure the precedence between the service train and the actual train.
-In this way our model shares features of both microscopic and macroscopic models.
-At the decision stations it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting.
-Meanwhile from the point of view of the rest of the network it is a macroscopic model.
+Rolling stock circulation conditions are prescribed to ensure the precedence between the service train and the actual train.
+Our model shares features of both microscopic and macroscopic models.
+At the decision stations, it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting.
+Meanwhile, from the point of view of the rest of the network, it is a macroscopic model.
 4 Methods and Model
 In the following, we describe our model in detail. Section 4.1 defines sets, parameters, and decision variables. Section 4.2 describes our Integer Linear Programming (ILP) formulation.
-4.1 Sets, parameters and decisions
-To formulate our decision variables, constraints, and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan.
+4.1 Sets, parameters and decision variables
+In order to define our decision variables, constraints, and objective function, we first determine sets of index tuples from the given infrastructure data, timetable data, and the rolling stock circulation plan.
 Also, we introduce parameters calculated from the same input.
-Sets
-Let us denote by $\hat{S}$ the set of all stations, and by $S\subset\hat{S}$ the decision stations we have determined in advance, as already described.
+4.1.1 Sets
+$\hat{S}$ stations
+$S$ decision stations
+$J$ trains
+$\mathcal{S}_{j}$ stops of train $j$ (ordered)
+$\mathcal{J}^{2\,(\text{close})}$ train pairs with precedence variables
+$\mathcal{J}^{2\,(\text{single})}$ train pairs possibly conflicting on single-track segments
+$\mathcal{J}^{2\,(\text{headway})}$ train pairs prone to violation of minimal headway
+$\mathcal{J}^{2\,(\text{turn})}_{s}$ train pairs with the same rolling stock turning at $s$
+$\mathcal{J}^{2\,(\text{track})}_{s}$ train pairs possibly competing for the same track of $s$
+$\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}$ train pairs possibly competing for an interlocking area of $s$ upon departure
+$\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon the departure of $j$ from $s$ while $j^{\prime}$ arrives to $s$ from the direction of $s^{\prime}$.
+$\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival from $s^{\prime}$ under special circumstances described in the main text.
+$\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival, under special circumstances described in the main text.
+$\mathcal{C}^{2}_{j}$ station pairs subsequent in the route of $j$
+$\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ station pairs subsequent in the route of both $j$ and $j^{\prime}$ heading in the same direction
+$\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ station pairs which are connected with a single-track line segment, subsequent in the route of both $j$ and $j^{\prime}$ heading in the opposite directions
+Table 1: Summary of sets
+Let $\hat{S}$ denote the set of all stations, and $S\subset\hat{S}$ the decision stations which have been selected in advance, as already described.
 The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path.
-Set $\mathcal{S}_{j}$ includes decision stations only and it is an ordered set.
+The sets $\mathcal{S}_{j}$ list decision stations only and they are ordered.
 In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following.
-Set $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J}\left(\forall s\in\mathcal{S}\right)$ is the set of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair.
-This set is deduced from the rolling stock circulation plan and the timetable.
-Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains that are close enough to each other in time so that precedence variables have to be defined for them.
-More details will be given in the description of the parameter $d_{\max}$ which this set depends on.
-Set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that can potentially violate the minimal headway time between two subsequent trains on a line segment.
-For a pair of trains to be in $\mathcal{J}^{2\,(\text{headway})}$, both their routes need to include the same line segment so that the trains are moving in the same direction on it and can meet there according to model parameters, i.e.
-maximal allowed delay.
-Set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of all trains that share at least one single-track line segment as a common part of their routes so that they are heading in the opposite direction.
-Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to occupy the same track on station $s$ anytime during the planning horizon, thereby competing for the same track.
-Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their departure from station $s$.
-Set $\left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ in order to have $j$ to departure $s$ while $j^{\prime}$ arrive $s$ from the direction of $s^{\prime}$.
-Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-P possibility for them between $s$ and $s^{\prime}$ Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ so that either they both come from the direction of $s^{\prime}$ but there is a M-P possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$.
-Set $\left(\forall j\in\mathcal{J}\right)\ \mathcal{C}^{2}_{j}$ is the set of all station pairs that are subsequent in the route of $j$.
-Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction.
-Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions.
+  • The set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of train pairs for which a precedence variable can potentially be defined in the model.
+It could be considered to be the whole $\mathcal{J}\times\mathcal{J}$, but this would result in a significant number of redundant variables.
+At the description of the parameter $d_{\max}$ we will introduce a way how to initially reduce the number of train pairs potentially involved in any constraint; essentially this will be the set of trains that potentially “interact”.
+This set will be divided into subsets related to specific “interactions” in what follows.
+  • The set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ contains all train pairs that can potentially meet and pass on a single-track line segment.
+This set can be obtained by considering all pairs of trains that are heading in the opposite direction on any single-track line, assuming all possible delays.
+If their train paths possibly intersect on any single-track segment, they are in $\mathcal{J}^{2\,(\text{single})}$.
+  • The set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ contains all train pairs that can violate the minimal headway condition, that is, those which potentially follow each other heading in the same direction, assuming any of the possible delays that can occur in the model instance.
+  • The sets $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair.
+These sets are deduced from the rolling stock circulation plan and the timetable.
+  • The sets $\mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly occupy the same track on the station $s$, thereby competing for the same track.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their departure from station $s$.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$.
+For these train pairs, the respective constraints will ensure that they do not pass the same interlocking area at the same time upon departure.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$.
+The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same interlocking area at the same time upon arrival, which is also enforced by the interlocking system.
+The possibility for M-O between the current and previous decision stations (e.g.
+using multiple tracks) is to be modeled specifically, hence the distinction between these sets and the $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$-s.
+  • The sets $\mathcal{C}^{2}_{j},\ \left(\forall j\in\mathcal{J}\right)$ are the sets of all station pairs that are subsequent in the route of $j$.
+  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction.
+  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions.
 The order within the pairs is determined by $j$.
-All these sets can be enumerated based on the input data in a straightforward manner.
-Parameters
-The parameters that appear in our model are the following:
+All these sets can be enumerated based on the input data in a straightforward manner. A brief summary of the sets is provided in Table 1.
+4.1.2 Parameters
+$\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ running time of $j$ from $s$ to $s^{\prime}$
+$\tau^{\text{(headway)}}(j,j^{\prime},s\rightarrow s^{\prime})$ minimal headway time for $j^{\prime}$ following $j$ from $s$ to $s^{\prime}$
+$\tau^{(\text{dwell})}(s,j)$ minimal dwell time of $j$ at $s$
+$\tau^{\text{(turn)}}(s,j,j^{\prime})$ rolling stock turnaround time for $j$ to continue as $j^{\prime}$ from s
+$\sigma(j,s)$ originally scheduled departure of $j$ from $s$
+$\upsilon(j,s)$ earliest possible factual departure time of $j$ from $s$
+$d_{\text{max}}$ upper bound for secondary delays
+Table 2: Summary of parameters
+The parameters that appear in our model, also summarized in Table 2, are the following:
   • $\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ is the running time of $j$ from $s$ to $s^{\prime}$.
   • $\tau^{\text{(headway)}}(j,j^{\prime},s\rightarrow s^{\prime})$ is the minimal headway time for $j^{\prime}$ following $j$ from $s$ to $s^{\prime}$.
-  • $\tau^{(\text{switch})}(j,j^{\prime},s)$ is the running time of $j$ over a interlocking area of $s$, where $j$ may be in conflict with $j^{\prime}$.
-(In our examples we do not consider cases when, e.g., on bigger stations there are multiple switches with different technological times.
-The model could trivially be extended to cover such scenarios by adding extra indices if relevant.)
+  • $\tau^{(\text{switch})}(j,j^{\prime},s)$ the time for which train $j$ occupies an interlocking area at $s$; this time should elapse before $j^{\prime}$ can start occupying the same area.
+Our model is simplified in this respect: we consider an upper estimate of this time valid for any switching zone of the station.
+The assumed values are determined from the parameters of station technology and the characteristics of the train pairs.
   • $\tau^{(\text{dwell})}(s,j)$ is the minimal dwell time of $j$ at $s$.
   • $\tau^{\text{(turn)}}(s,j,j^{\prime})$ is the minimum turnaround time for the rolling stock of a train $j$ terminating at $s$ to continue its journey as train $j^{\prime}$.
-  • $\sigma(j,s)$ is the scheduled departure time of $s$ from $j$, given by the original timetable.
+  • $\sigma(j,s)$ is the scheduled departure time of $s$ from $j$, according to the original timetable.
   • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance in the network happens.
-It is the maximum of $\sigma(j,s)$ and the technically feasible earliest departure time, and may be more constraining than the first depending on the initial delays.
-The latter is calculated assuming that the train follows its planned route at minimum running time and that there are no other trains present on the network.
+It is the maximum of the planned departure time $\sigma(j,s)$, and the technically feasible earliest departure time.
+The latter is calculated using the imaginary situation that the train starts from the initial location at the initial time defined by the initial delays (from the problem instance), and it moves to station $s$ as fast as technically possible, without having any other trains on the network, and then departs from there after the minimal stopping time has elapsed.
+These parameters are calculated before the optimization for the problem instances, based on the technological data of the network (minimum passing times) and the initial delays.
   • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$.
-This parameter sets an upper limit for the possible secondary delays that can arise on a particular part of the network.
-Setting such a bound is common in the literature, see e.g.
-the work of D’Ariano et al.
-(2007).
+This parameter sets an upper limit for the possible secondary delays.
+Setting such a bound is not uncommon in the literature, see e.g.
+the work of [10].
 We set the same value of this parameter for all our testing instances.
@@
 Meanwhile, it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
-The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $J^{2\,(\text{close})}$.
-A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$.
-The number of constraints will be proportional to the average number of trains that can meet another train at a station; this will be linear in $d_{\max}$.
-Decision variables
-Our decision variables are the following. First, we use departure time variables:
-$t^{(\text{out})}(j,s)\in\mathbb{N}$ (1)
+As mentioned in Section 4.1.1, the parameter $d_{\max}$ makes it possible to define $\mathcal{J}^{2\,(\text{close})}$, the set of trains potentially appearing in constraints, given a problem instance.
+Namely, a pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet at any station, given the original timetable, the initial delays, and assuming that no train can have a secondary delay greater than $d_{\max}$.
+The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $\mathcal{J}^{2\,(\text{close})}$ with smaller cardinality.
+Choosing too small $d_{\max}$ can result in the infeasibility of the model.
+4.1.3 Decision variables
+$t^{(\text{out})}(j,s)$ departure time of train $j$ from decision station $s$
+$y^{(\text{out})}(j,j^{\prime},s)$ order of trains $j$ and $j^{\prime}$ upon their departure from $s$
+$y^{(\text{in})}(j,j^{\prime},s)$ order of trains $j$ and $j^{\prime}$ upon their arrival from $s$
+$z(j,j^{\prime},s,s^{\prime})$ the order of trains at a resource such as e.g. a single-track segment, located between $s$ and $s^{\prime}$ the trains $j$ and $j^{\prime}$ are competing for.
+Table 3: Summary of decision variables
+Our decision variables, summarized in Table 3, are the following. First, we use departure time variables:
+$t^{(\text{out})}(j,s)\in\mathbb{N},\ \ \ \text{where}\ \ \ \upsilon(j,s)\leq t^{(\text{out})}(j,s)\leq\upsilon(j,s)+d_{\max},$ (1)
 defining the departure time of train $j\in\mathcal{J}$ from station $s\in\mathcal{S}_{j}$.
 Such a variable is defined for all decision stations $\mathcal{S}_{j}$.
-In the case of trains that terminate within the modeled part of the network, the last station is not included in $\mathcal{S}_{j}$ The arrival time of the trains at stations, $t^{(\text{in})}(j,s)\in\mathbb{N}$ are trivially related to the $t^{(\text{out})}(j,s)$ variables through a constant offset (c.f.
-Eq.
-(3)), given the fixed running time assumption.
-Second, in addition to the time variables, we use three sets of binary precedence variables.
-Decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: the variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and 0 otherwise.
+The bounds follow trivially from the notion of the $\upsilon(j,s)$ and $d_{\text{max}}$ parameters.
+In the case of trains that terminate within the modeled part of the network, the last station is not included in $\mathcal{S}_{j}$.
+For the sake of notational convenience, we also introduce the arrival time of the trains at stations, $t^{(\text{in})}(j,s^{\prime})\in\mathbb{N}$.
+Having assumed fixed running times (green way policy), arrival times are defined by the previous departure times, (c.f.
+Equation (3)).
+Hence, arrival times are not independent decision variables themselves.
+This explains why we need sets and decision variables related to various combinations of inbound and outbound trains when describing conflicts in switching zones.
+In addition to the time variables, we use three sets of binary precedence variables.
+The decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: such a variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and it is 0 otherwise.
 The $(j,j^{\prime},s)$ tuples for which an $y$ variable is defined will be specified later.
 Similarly, the precedence variables $y^{(\text{in})}(j,j^{\prime},s)\in\{0,1\}$ prescribe the order at the entry to stations; the value is $1$ if $j$ arrives to $s$ before $j^{\prime}$.
-The last set of precedence variables describes the precedence of trains at some resource located between stations $s$ and $s^{\prime}$ (e.g.
-a single-track segment used by $j$ and $j^{\prime}$ heading in opposite direction).
+The last set of precedence variables describes the precedence of trains $j$ and $j^{\prime}$ competing for a resource located between stations $s$ and $s^{\prime}$.
 The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$.
-Also in this case, the quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later.
+The quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later, but let us provide an example here.
+Consider two stations $s$ and $s^{\prime}$ that are connected by a single-track line segment, and trains $j$ and $j^{\prime}$ are heading in opposite directions, competing for the single-track segment as a resource.
+In this case, the variable describes which of the trains can occupy the single-track link first.
+(Recall that M-P is possible on decision stations only.)
 4.2 ILP formulation
 Given the index sets, variables, and the parameters and decision variables of the model, now we formulate the optimization objective and the constraints.
-Objective function
-Our goal is to minimize the secondary delays that occur in the analyzed part of the railway network.
-Hence, a suitable objective function is the weighted sum of secondary delays at the destination station:
-$f(t)=\frac{1}{d_{\text{max}}}\sum_{j\in\mathcal{J}}w(j)\left(t^{(\text{out})}(j,s^{*})-\upsilon(j,s^{*})\right).$ (2)
+4.2.1 Objective function
+Our goal is to minimize the secondary delays that occur in the analyzed part of the railway network. A suitable objective function is the weighted sum of secondary delays at the destination station:
+$f(t)=\sum_{j\in\mathcal{J}}w_{j}\left(t^{(\text{out})}(j,s^{*})-\upsilon(j,s^{*})\right),$ (2)
 where $s^{*}$ is the last element of $S_{j}$, and $w_{j}$ are weights for each train representing its priority.
-The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as the weighted sum of secondary delays on all stations is straightforward.
-The constant multiplier $1/d_{\text{max}}$ is optional; we use it for better comparability of different instances.
-Constraints
+The weights have to be chosen in such a way that they reflect the rules and practices of the local dispatching system.
+As these rules are often formulated only qualitatively, the particular values are somewhat arbitrary; the model can be fine-tuned in this respect.
+(In Section 6.1 we provide more details of the weight assignments for Polish PKP PLK.) The reason for restricting ourselves to the secondary delays at the destination stations only is that this is the actual figure of merit chosen by the respective railway authorities; the generalization to other linear objectives such as, e.g., the weighted sum of secondary delays on all stations, is straightforward.
+4.2.2 Constraints
 The constraints of the model are the following.
@@
 Each train needs a minimal time to get to the subsequent station:
-$\forall j\in\mathcal{J}\ \forall(s,s^{\prime})\in\mathcal{C}_{j}\quad t^{(\text{in})}(j,s^{\prime})=t^{(\text{out})}(j,s)+\tau^{(\text{pass})}(j,s\rightarrow s^{\prime})$ (3)
+$t^{(\text{in})}(j,s^{\prime})=t^{(\text{out})}(j,s)+\tau^{(\text{pass})}(j,s\rightarrow s^{\prime}),\quad\forall j\in\mathcal{J},\ \forall(s,s^{\prime})\in\mathcal{C}_{j}.$ (3)
+Recall that this condition represents the green way policy and makes the variables $t^{(\text{in})}$ depend trivially on the $t^{(\text{out})}$-s.
+The reason for keeping $t^{(\text{in})}$-s is that they make the model description more transparent.
 Headways
 A minimal headway time is required between subsequent train pairs on the common part of their route as
-$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}t^{(\text{out})}(j^{\prime},s)\geq$
-$t^{(\text{out})}(j,s)+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})-C\cdot y^{(\text{out})}(j^{\prime},j,s),$ (4)
-where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$.
-(Constraints of this structure are often termed as big-M constraints in this context.) In our implementation, we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
-$C=-\min(t^{(\text{out})}(j^{\prime},s))+\max(t^{(\text{out})}(j,s))+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})$ (5)
+$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime})-C\cdot y^{(\text{out})}(j^{\prime},j,s),$
+$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})},\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},$ (4)
+where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$: in this case, $j^{\prime}$ goes first, so the constraint should be trivially satisfied.
+Otherwise speaking, the constraint should only be active for $y^{(\text{out})}(j^{\prime},j,s)=0$.
+(Constraints of this structure are often termed as big-M constraints in this context.) In order to find the smallest value of $C$ for this to hold, we can exploit the bounds on the time variables in Equation (1)): we have
+$t^{(\text{out})}(j^{\prime},s)\geq\upsilon(j^{\prime},s)$ (5)
+and
+$t^{(\text{out})}(j,s)\leq\upsilon(j,s)+d_{\text{max}},$ (6)
+hence, setting
+$C=-\upsilon(j^{\prime},s)+\upsilon(j,s)+d_{\max}+\tau^{(\text{headway})}(j,j^{\prime},s\rightarrow s^{\prime}),$ (7)
+if $y^{(\text{out})}(j^{\prime},j,s)=1$ then Equation (4) will take the form
+$t^{(\text{out})}(j^{\prime},s)-\upsilon(j^{\prime},s)\geq t^{(\text{out})}(j,s)-d_{\text{max}}-\upsilon(j,s),$ (8)
+which always holds as desired. Therefore, $C$ will be set according to Equation (7).
 Single-track occupancy
 Trains moving in opposite directions cannot meet on the same single-track line segment:
-$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{single})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$
-$t^{(\text{out})}(j^{\prime},s^{\prime})\geq t^{\text{(in)}}(j,s^{\prime})-C\cdot z(j^{\prime},j,s^{\prime},s),$ (6)
-where $C$ is a big enough constant chosen similarly to that in Eq. (4).
+$t^{(\text{out})}(j^{\prime},s^{\prime})\geq t^{\text{(in)}}(j,s^{\prime})-C\cdot z(j^{\prime},j,s^{\prime},s),$
+$\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{single})},\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},$ (9)
+where $C$ is a big enough constant chosen similarly to that in Equation (4).
 Minimal dwell time
 Each train has to occupy the station node for a prescribed time duration at each station:
-$\forall j\in\mathcal{J}\forall s\in S_{j}\quad t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s).\$ (7)
-Timetable
-No train is allowed to depart before its scheduled departure time:
-$\forall j\in\mathcal{J}\forall s\in S_{j}\quad t^{(\text{out})}(j,s)\geq\sigma(j,s).$ (8)
+$t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s),\ \forall j\in\mathcal{J},\ \forall s\in S_{j}.$ (10)
 Station track occupancy
 Station tracks can be occupied by at most one train at a time:
-${\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
-where $C$ is chosen similarly to that in Eq.
-(4) again.
+$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\ {\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}},\$ (11)
+where $C$ is chosen similarly to that in Equation (4) again.
 Note that this requirement may not be needed for depot tracks; the exceptions can be handled by the proper definition of $\mathcal{J}^{2\,(\text{track})}_{s}$.
@@
 These ensure that trains cannot meet in interlocking areas:
-$\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}t^{(\text{out})}(j^{\prime},s)\geq$
-$t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$ (10)
-$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
-$t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot z(j^{\prime},j,s^{\prime},s),$ (11)
-$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
-$t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s^{\prime}),$ (12)
-$\forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}t^{(\text{in})}(j^{\prime},s)\geq$
-$t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{in})}(j^{\prime},j,s),$ (13)
-with a choice of $C$ similar again to that in Eq. (4).
+$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$
+$\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s};$ (12)
+$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot z(j^{\prime},j,s^{\prime},s),$
+$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}};$ (13)
+$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s^{\prime}),$
+$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}};$ (14)
+$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{in})}(j^{\prime},j,s),$
+$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}},$ (15)
+with a choice of $C$ similar again to that in Equation (4). These constraints are also necessary for to make the realization of the modified timetable possible using the interlocking systems.
 Rolling stock circulation constraints
-These are introduced to bind the train with the shunting movement called also service train.
+These are introduced to bind a train with a shunting movement, termed also as a service train.
 If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
-$\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s).$ (14)
+$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s),\ \forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}.$ (16)
 Order of trains
 We have additional conditions on the $y$-variables, concerning the case when M-O is not possible on the line or station.
-$\begin{split}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}\\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\\ y^{(out)}(j,j^{\prime},s)=y^{(out)}(j,j^{\prime},s^{\prime})\\ \forall(s,s^{\prime})\in S^{\times 2}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\\ y^{(in)}(j,j^{\prime},s^{\prime})=y^{(out)}(j,j^{\prime},s^{\prime})\end{split}$ (15)
-The objective function in Eq. (2), together with the constraints in Eq. (3)-(14) define our mathematical programming model.
-The model yields an integer linear program, and the time variables can be constrained even into a finite range using the parameter $d_{\max}$.
-Note that the binary variables have the obvious symmetry property
-$\begin{split}y^{(\text{out})}(j^{\prime},j,s)=1-y^{(\text{out})}(j,j^{\prime},s)\\ z(j^{\prime},j,s^{\prime},s)=1-z(j,j^{\prime},s,s^{\prime}),\end{split}$ (16)
-which is taken into account explicitly by using independent binary variables only.
-As for the scaling of our model, the number of time variables i.e.
-#$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  Domino et al.
-(2022b)).
-To obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations.
-We assume that in average each train visits $\alpha\in(0,1)$ fraction of the stations, we have then
-#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}.$ (17)
+In particular, assume that the trains $j$ and $j^{\prime}$ are following each other from $s$ to $s^{\prime}$, heading in the same direction.
+If they use the same track on $s^{\prime}$, they will have to leave $s^{\prime}$ in the same order as they left $s$.
+Formally:
+$\begin{split}y^{(\text{out})}(j,j^{\prime},s)=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(j_{0},j_{0}^{\prime})\in\mathcal{J}^{2\,(\text{headway})},\ \forall(s,s^{\prime})\in\mathcal{C}^{2\,(\text{common})}_{j_{0},j_{0}^{\prime}}\\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{headway})}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}\end{split}$ (17)
+Similarly, if trains $j$ and $j^{\prime}$, enter $s^{\prime}$ from any location, they are passing the switching zone of station $s^{\prime}$, then they are passing the station $s^{\prime}$ in the same direction using the same track, they have to keep their order upon departure.
+Formally:
+$\begin{split}y^{(\text{in})}(j,j^{\prime},s^{\prime})=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(s,s^{\prime})\in S\times S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{\text{in}},\text{MO})}_{s^{\prime},s}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}.\end{split}$ (18)
+The objective function in Equation (2), together with the constraints in Equation (3)-(16) define our mathematical programming model.
+Our model is formulated as an integer linear program.
+Moreover, as the time variables are constrained a finite range using the parameter $d_{\max}$, all variables are of a finite range.
+We intend to limit the number of variables as much as possible.
+The binary variables have the obvious symmetry property
+$\begin{split}y^{(\text{out})}(j^{\prime},j,s)=1-y^{(\text{out})}(j,j^{\prime},s)\\ z(j^{\prime},j,s^{\prime},s)=1-z(j,j^{\prime},s,s^{\prime}).\end{split}$ (19)
+These are added to the model as constraints, in a way that only half of the variables are considered as independent, and the rest are removed.
+4.2.3 Scaling analysis
+Let us discuss the number of variables and constraints in our model.
+The number of time variables #$\#t$, the product #$\#\mathcal{J}\#\mathcal{S}$ of the number of trains and the number of stations serves as a trivial but rough upper bound.
+(See also Section 3.1 of the work of  [17]).
+To obtain a more accurate estimate of the number of these variables we observe that in our model no train visits all the stations.
+To provide an estimate, we introduce $\alpha\in(0,1)$ as the fraction of the stations visited by an average train.
+Thus we have
+#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}\leq\#\mathcal{J}\#\mathcal{S}.$ (20)
 as a tighter bound.
-Suppose, that most of the network is composed mainly of double-track lines.
-Then there will be typically a single precedence variable per train and station ($y^{(out)}(j,j^{\prime},s)$), and thus their number can be estimated by:
-#$\#y+\#z\approx\#y\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S},$ (18)
-where $N(d_{\max})$ is the number of trains the train can meet at a station on average. Under such assumptions, the total number of variables grows linearly with number of trains and stations:
-#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}\#\mathcal{S}.$ (19)
-If, on the other hand, the network has dominantly single-track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
-#$\#y=\#z\approx\alpha N(d_{\max})\#\mathcal{J}\#\mathcal{S}.$ (20)
-In both cases we assume that there are not many ($y^{(in)}(j,j^{\prime},s)$) variables, as they are tied certain particular interlocking area conditions.
-Based on Eq.
-(16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single-track occupation constraints.
-For the interlocking area occupation constraints, it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable.
-For the running time and minimal dwell time we expect one constraint per train and station.
-This number of constraints can be estimated by
-#$\#constr.\approx 6\#y+2\#z+2\#\mathcal{J}\#\mathcal{S}$ (21)
-where the second term is tied to single-track line conditions (one condition per $z$ variable).
-#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double-track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single-track}.\end{cases}$ (22)
-Because of the limitations in time imposed by $d_{\max}$ the model can be viewed as linear in number of trains and number of stations.
-Hence, as expected, the model is more complex for the network in which single-track lines dominate.
-Although the analyzed network is mostly composed of the double-track lines, we will analyze certain use-cases with the fraction of single-track lines increased.
-Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables, and that of the number of constraints in the number of trains.
+In what follows we will mostly $\alpha=1$, the rough bound, however, we will keep the parameter $\alpha$ in the formulae to reflect its role in the scaling.
+Let us assume, that half the of trains (#$\frac{\#\mathcal{J}}{2}$) are going in one direction and the other half are going in the other.
+Then the number of trains each one meets going in the same direction is $\alpha\frac{\#\mathcal{S}\#\mathcal{J}}{2}$.
+This yields #$\#\mathcal{J}\cdot\frac{\alpha\#\mathcal{S}\#\mathcal{J}}{2}$ $y^{(\text{out})}$ variables.
+However, because of the symmetry property in Equation (19), only half of these variables are independent, thus we have $\alpha\#\mathcal{S}\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{out})}$ of these variables.
+As for the trains going in opposite directions, we assume that each train meets at most #$\frac{\#\mathcal{J}}{2}$ trains on each station in the single-track parts of the network.
+Hence, the number of $z$ variables is at most #$\#\mathcal{J}\cdot\frac{1}{2}\cdot\frac{\alpha\#\mathcal{J}(\#\mathcal{S}-1)}{2}$ (here, #$\#\mathcal{S}-1$ comes from the observation that the $z$-variables are tied to pairs of stations).
+Variables $y^{(\text{in})}$ occur mainly in Eqs.
+(4.2.2)-(4.2.2), relevant for a few stations only: there is typically one such station in the network, yielding at most #$\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{in})}$ variables.
+In summary, the number of precedence variables can be estimated as
+#$\begin{split}&\#y^{(\text{out})}\leq\alpha\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\leq\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\\ &\#y^{(\text{in})}\leq\frac{\#\mathcal{J}^{2}}{4}\\ &\#z\leq\alpha\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\leq\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\\ &\#y+\#z=\#y^{(\text{out})}+\#y^{(\text{in})}+\#z<\frac{\#\mathcal{J}^{2}}{2}\#\mathcal{S}\end{split}$ (21)
+Note that these bounds are rather pessimistic: setting the parameter $d_{\max}$ as an upper bound on the secondary delays implies a reduction in the number of trains that can potentially conflict.
+However, the actual number of them can depend on the train density and structure of the timetable in various ways, hence, it is not possible to provide a tighter bound that would be valid in general.
+Finally, let us remark as the $z$-variables are mainly tied to single-track traffic, the model is more complex for networks dominated by single-track segments.
+Based on Eqs. (20) and Equation (21), the scaling of the upper bound on the the total number of decision variables can be estimated by
+#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}^{2}\#\mathcal{S}+\text{const'}\#\mathcal{J}\#\mathcal{S},$ (22)
+which is quadratic in the number of trains and linear in number of stations.
+Now let us estimate the number of constraints.
+Considering the headway constraints in Equation (4), there is one such constraint for each variable.
+However, because of the symmetry property in Equation (19), only half of the variables are independent, hence, the number of these constraints is twice the number of independent variables.
+Along the same arguments, in case of the constraints in Equation (11), i.e.
+station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each $z$ variable.
+It is more complicated to estimate the number of interlocking area occupation constraints in Eqs.
+(4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables.
+A good assumption is to consider $2$ constraints for each $y^{(\text{out})}$ variable, $2$ constraints for each $y^{(\text{in})}$ variable, and $2$ constraints for each $z$ variable.
+Further, we expect one minimal dwell time constraint per train and station (described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term).
+Concerning train order constraints in Eqs (17) (18), we also expect one constraint per train and station (also described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term).
+We exclude the rolling stock circulation constraints from this consideration as their contribution to the number of constraints is very limited.
+In summary, the total number of constraints obeys the following upper bound:
+#$\#constr.<6\#y^{(\text{out})}+2\#y^{(\text{in})}+4\#z+2\alpha\#\mathcal{J}\#\mathcal{S}.$ (23)
+Substituting the estimated number of variables in Equation (21) we have
+#$\#constr.<\alpha\frac{3}{2}\#\mathcal{J}^{2}\#\mathcal{S}+\frac{1}{2}\#\mathcal{J}^{2}+\alpha\#\mathcal{J}^{2}(\#\mathcal{S}-1)+2\alpha\#\mathcal{J}\#\mathcal{S}$ (24)
+Setting $\alpha=1$ as the worst case, we get
+#$\#constr.<\frac{5}{2}\#\mathcal{J}^{2}\#\mathcal{S}-\frac{1}{2}\#\mathcal{J}^{2}+2\#\mathcal{J}\#\mathcal{S}.$ (25)
+Let us stress that the provided generic pessimistic upper bounds only serve the purpose of getting an overall picture of the scaling; the particular number of variables and constraints is lower and will be specified for our problem instances.
 5 Hybrid quantum-based approach
-To overcome the limitations described in in Section 2 of current hardware quantum annealers, we apply a hybrid (quantum-classical) solver.
-In particular, we use the ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a); a cloud-based proprietary solver that is developed by the market leader of QA hardware and is available as a service.
-It utilizes a hybrid approach that combines classical computational power with quantum processing.
-In particular, we have used the Constrained Quadratic Model (CQM) (D-Wave Quantum Inc., 2022b) Solver. The CQM hybrid solver works as follows.
+To overcome the limitations of current hardware quantum annealers described in Section 2, we apply a hybrid (quantum-classical) solver.
+In particular, we use the ’Leap Hybrid Solver Service (HSS)’ [8]; a cloud-based proprietary solver that is developed by the market leader of QA hardware and is available as a service.
+It implements a hybrid approach that combines classical computational power with quantum processing.
+In particular, we have used the Constrained Quadratic Model (CQM) [9] Solver.
+The CQM hybrid solver works as follows.
 Input
@@
 Preprocessing
-Preprocessing includes subproblem identification and decomposition to smaller instances (Tran et al., 2016).
+Preprocessing includes subproblem identification and decomposition to smaller instances [38].
 Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
@@
 In the course of the solution process the hardware quantum annealer is also invoked, to solve smaller QUBO problems that can potentially boost the classical heuristics.
-Integer problems are transformed to binary using state-of-the-art methods like the one described by Glover et al.
-(2019); Karimi and Ronagh (2019), generally applicable for integer problems.
+Integer problems are transformed to binary using state-of-the-art methods like the one described by [19, 24], generally applicable for integer problems.
 When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions.
@@
 In this way, even though the physical annealer supports problems with limited size (e.g.
-the Pegasus machine has $5500$ physical qubits (Dattani et al., 2019), and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process.
-The final solution is composed by the solver’s module.
+the Pegasus machine has $5500$ physical qubits [12], and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process.
+The final solution is assembled by the solver’s module.
 Output
@@
 The CQM (integer) solver returns a series of solutions, which can be an additional asset.
-Nevertheless, for sake of presentation, our approach is to analyze only the best CQM solution given each run of the hybrid algorithm (realization).
+For the probabilistic nature of the solver, the returned set of solutions is often termed as a sample, and the process of obtaining the sample is termed as sampling.
+Note that CQM, like most heuristic solvers, does not warrant feasibility by default, hence, a direct check is performed after obtaining the possible solutions.
+Nevertheless, for the sake of presentation, our approach is to analyze only the best CQM solution given each run of the hybrid algorithm (realization).
 While D-Wave’s proprietary solvers act as black boxes, i.e.
@@
 As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min.
-This is a time limit for heuristics running in parallel(D-Wave Quantum Inc., 2022b): unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution.
+This is a time limit for heuristics running in parallel[9]: unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution.
 We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting on the model performance and solution quality for our problem.
-The CQM Solver can solve problems encoded in the form of Eq.
-(2) – (15) (c.f.
+The CQM Solver can solve problems encoded in the form of Equation (2) – (19) (c.f.
 Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints.
 Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum-classical) solvers.
-As a classical solver, IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) was used.
+As a classical solver, we have opted for IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) as one of the most prevalently used commercial solver.
 The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory.
 As a hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.
+In addition to solving the problems to optimality, we include CPLEX results with limited computational time, for the sake of a fair comparison with heuristics.
 6 Computational results
-In this section, we demonstrate the performance of the proposed hybrid (quantum-classical) approach on a network at Polish railways - the central part of the Upper Silesia Metropolis.
+In this section, we demonstrate the performance of the proposed hybrid (quantum-classical) approach on a network of Polish railways - the central part of the Upper Silesia Metropolis.
 In particular, we compare the performance of D-Wave’s hybrid quantum-classical solver (CQM) with CPLEX.
-The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances including various initial delays and track closures.
-As the objective is tied to secondary delays, we can roughly apply it to assess the degree of difficulty of the problem.
+The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different levels of disturbances including various initial delays and track closures.
+The considered closures are situations in which one or more tracks between decision stations are closed and the traffic has to be redirected to the remaining tracks between decision stations.
+This can imply the need to add extra decision stations to the model to bind the part of the line with the closure; this also occurs in our examples.
 Section 6.1 describes the considered network and traffic characteristics.
-Section 6.2 presents results concerning generic examples on the core line of the network, comparing double-track-line scenarios with single-track-line scenarios and closure scenarios.
-Section 6.3 presents analogous results for larger network with real-life timetable.
+Section 6.2 presents results concerning generic synthetic examples on the core line of the network, comparing double-track-line scenarios with single-track-line scenarios and closure scenarios.
+Section 6.3 presents similar results for larger network with real-life timetable.
 Here the case studies concern various scenarios with different degrees of difficulty.
@@
 Polish State Railways, Polish Railway Lines).
-Figure 2: Selected part of Polish railway network located in the central part of the Metropolis and examined in our research.
+Figure 2: The part of the Polish railway network which is the subject of our considerations.
+It is located in the central part of the Metropolis.
+The displayed non-decision stations such as ZZ or RCB become decision stations in certain problem instances, e.g.
+due to an assumed closure of one of the tracks between ZZ and RCB.
 The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analyzed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single-track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O).
-As described before, the decision stations are the only ones that appear explicitly in our model.
-Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made with respect to the decision stations.
-All other stations and junctions are treated as line blocks as long as no rerouting or retracking is considered; they will appear in the model implicitly via parameters.
+As described before, the decision stations are the only ones that appear explicitly in our model by assigning decision variables.
+Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made concerning the decision stations.
+All other stations and junctions are treated as line blocks as long as no rerouting or retracking (e.g.
+in the case of closure) is considered; they will appear in the model implicitly via parameters.
+In such a case precedence of trains cannot be changed on these stations as there are no decision variables assigned to them.
 These stations include:
@@
 […]
-In all computation priority weights in Eq.
-(2) we use $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable).
+The priority weights in Equation (2) are set to $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable), in all the computations.
+These weights result from interpreting the instructions on train traffic management rules and priority levels for trains set out in the ”Instruction for train operation (Ir-1)” 222see the paragraph $51$ of Instrukcja o prowadzeniu ruchu pociagow Ir-1 (in Polish), Eng.
+Instruction for railway traffic https://www.plk-sa.pl/files/public/user_upload/pdf/Akty_prawne_i_przepisy/Instrukcje/Wydruk/Instrukcja_o_prowadzeniu_ruchu_pociagow_Ir-1_-_wchodzi_w_zycie_od_14_sty....pdf accessed 18 November 2024 and the real-life dispatching practices on the PKP PLK network.
+The default value for regular traffic is $w_{j}=1$, while there is no priority for shunting.
+Observe the higher priorities of InterCity trains and the even slightly higher priorities of express trains operated by Express Intercity Premium bullet trains which are briefly referred to as “express trains”.
+These intercity trains and express trains have strong priority in the case of small-scale disturbances, but this priority can be neglected in the case of larger disturbances.
 As for the solver parameter t_min, after initial experiments, we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality.
@@
   1. 1.
-    $Line1$, the double-track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
+    $line1$, the double-track line with dense traffic (As illustrated in Fig 2, we have $3$ decision stations: KO, CB and GLC).
+We use 3 periods of an hourly repeating timetable with $10$ trains each hour and in each direction.
+Here, and also in the following lines we did not consider one of the trains for practical reasons, thus we have $59$ trains in a three-hour time horizon.
   2. 2.
-    $Line2$, similar to $line1$, with the additional closure of one of the tracks between ZZ and RCB.
-We use a cyclic $2$-hour timetable with $10$ trains each hour and each direction, i.e.
-we have $40$ trains.
+    $line2$, similar to $line1$, with the additional closure of one of the tracks between ZZ and RCB, hence, these stations are added to the set of decision stations, and single-track traffic is assumed between them.
+We have $39$ trains on a two-hour time horizon.
   3. 3.
-    $Line3$, the whole line is considered as a single-track, with a timetable of $3$ hours traffic and $21$ trains.
-For $line1$ and $line3$, the original timetable is feasible, while for $line2$ it is not feasible due to the closure.
+    $line3$, the whole line is considered a single track.
+As in the "$line2$" case, we add two extra decision stations, ZZ and RCB where M-P and M-O are possible.
+We have $21$ trains in a three-hour time horizon.
+For $line1$ and $line3$, the original timetable would be feasible without the initial delays, while for $line2$ it is not feasible due to the closure. The instances are summarized in Tab 4.
+line #$\#J$ n.o. hours decision stations track layout
+1 $59$ $3$ KO, CB, GLC double
+2 $36$ $2$ KO, CB, RCB, ZZ, GLC
+single between RCB and ZZ
+then double
+3 $21$ $3$ KO, CB, RCB, ZZ, GLC single
+Table 4: Features of synthetic instances.
 For each of the lines, we compute $12$ instances, each with different initial delays of train subsets.
-Instance 00 yields no initial delays.
+The initial delays were assigned manually, a-priory for each instance, choosing plausible values based on everyday practice.
+Instance 00 has no initial delays.
 Thereafter, we use the parameter value $d_{\max}=40\ \text{min}$, in all cases in this and subsequent subsection.
-Assuming that each train interacts with $N(d_{\max})=12$ trains for $line1$ and $line2$, and with $N(d_{\max})=6$ trains for $line3$, Eqs.
-(17) - (19), (22) allows us to quickly estimate the sizes of the problem.
-These sizes are compared with the computed (comp.) actual number of variables and constraints (mean over instances) in Tab.
-1.
-Here, $line2$ is the most complex in terms of constraints, which is reflected later by the computational difficulties of some of its instances.
-Furthermore, $line2$ is partially double-track and partially single-track, hence the actual number of constraints is bounded from above by the estimate for the single-track line and from below by the estimate for the double-track line.
-$Line1$ includes double-, whereas $line3$ single-track segments only.
-There actual number of constraints and variables appeared to be close to the estimate.
-l #$\#\mathcal{S}$$\#\mathcal{J}$ int vars # precedence vars # constraints
-comp. mean Eq. (17) act. mean
-Eq. (18)/Eq. (20)
-doub./sing.
-comp. mean
-Eq. (22)
-doub./sing.
-1 3 59 118 118 1538 1415 / - 7889 8731 / -
-2 5 40 142 133 1394 1600/3200 8491 6867/13067
-3 5 21 79 70 592 - / 840 3057 - / 3499
-Table 1: Synthetic experiments.
-The first column is the ordinal number of the line.
-For the estimated and actual number of variables and constraints, we use $\alpha=2/3$ in Eqs.
-(17) - (19), (22) (doub.
-- double-track, sing.
-- single-track, comp.
-mean - computed actual number of variables or constraints, mean over instances).
-The computational results are presented in Fig.
-3 a) b) c) for $line1$, $line2$, and $line3$, respectively.
-Each figure shows objective $\times$ $d_{\text{max}}$ (top part), computation time (comp.
+We chose its particular values by experimentation.
+The upper limits found in Section 4.2.3 are compared to the actual number of variables and constraints (mean over instances) in Table 5.
+The purpose of the table is to connect the theoretical discussion on scaling in Section 4.2.3 with the actual problem sizes.
+For example, $line2$ is the most complex in terms of constraints, which is reflected later by the computational difficulties of some of its instances.
+(Due to closure, $line2$ is partially double-track and partially single-track with heavy traffic.).
+The table serves also as the sanity check of our model.
+$line$$\#\mathcal{S}$$\#\mathcal{J}$ int vars # precedence vars # constraints
+actual mean
+Equation (20)
+upper lim
+actual mean
+Equation (21)
+upper lim
+acctual mean
+Equation (25)
+upper lim
+1 3 59 118 180 1538 5221 7889 24721
+2 5 39 142 200 1394 3802 8491 18642
+3 5 21 79 110 592 1102 3057 5502
+Table 5: Estimated and actual problem sizes on $line1$ $line2$ and $line3$ scenarios. For each scenario, the mean over instances was computed.
+The computational results are presented in Figure 3 a) b) c) for $line1$, $line2$, and $line3$, respectively.
+Each figure shows the objective (top part), computation time (comp.
 time, middle part), and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s.
 CPLEX results are also displayed.
-For scenario $line1$ (Fig.
-3 a) ), the CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$).
-Similarly, for $line3$ (Fig.
-3 c) ), only small deviations from the optimum can be observed for a number of instances.
-All computational times for $line1$ and $line3$ were significantly below $0.4$min, although CPLEX performed faster.
-Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objectives but t_min$=20s$ requires almost $4$ times longer computational time.
-For $line2$ (Fig.
-3 b) ), all CQM solutions were suboptimal, although feasible.
-More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than CPLEX, in several cases up to 5×5$\times$5 × faster.
+These latter include the optima as well as the heuristic results with computational time limited to at most 5 seconds.
+For the scenario $line1$ (Figure 3 a) ), the CQM solver found the optimal solutions in all of the cases except for instances $6$ and $11$.
+Similarly, for $line3$ (Figure 3 c) ), only small deviations from the optimum can be observed for a number of instances.
+All computational times for $line1$ and $line3$ were significantly below $30$seconds, although CPLEX was faster, and found the actual optimum before the possible 5 seconds limit.
+Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objective values, however, t_min$=20s$ requires almost $4$ times longer computational time.
+For $line2$ (Figure 3 b) ), all CQM solutions were suboptimal, although feasible.
+More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than exact CPLEX, in several cases up to $5$ times faster.
+CPLEX could not solve these instances to optimality when it was limited to 5 seconds; it returned close-to-optimal feasible solutions though.
 This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5).
-In principle, $line1$ (Fig.
-3 a) ) and $line3$ (Fig.
-3 c) ) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster.
-We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems.
-Figure 3: Synthetic experiments $Line1$, $Line2$, $Line3$ comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM.
+In principle, $line1$ (Figure 3 a) ) and $line3$ (Figure 3 c) ) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster.
+We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems, as the present instances for $line2$.
+The detailed analysis of the implications of suboptimality in CQM solutions for $line2$ from the railway point of view shows that there are two types of consequences: the increase of the passing time between subsequent nodes, and the change in certain trains’ order.
+Concerning the first issue, the realization of such suboptimal solutions can result in slowing down trains, which can be a motivation for further research, such as in the work of [14].
+Solutions affected by the second issue are suboptimal but still feasible and can be opted for by the human dispatcher because they have benefits for certain non-quantified reasons.
+Particular examples of real-life problem instances are presented in Appendix A.
+The hybrid quantum-classical solver CQM does not outperform purely classical heuristics in most problems at the present state of the art, and railway dispatching is not yet an exception.
+Our results demonstrate, however, that it already has the potential to handle difficult railway problems of moderate size.
+This implies that CQM, as a generic out-of-the-box solver, could already be a viable alternative to the currently existing solvers.
+Note that suboptimality can be accepted in many cases as it will be justified by the analysis of our real-world examples in Section 6.3.
+Further improvement in the runtime and the solution quality can possibly be achieved by developing custom classical-quantum heuristics tailored for the given problem, providing also information on when the quantum solvers can be the most efficient.
+Meanwhile, as there is a significant R&D effort put into the development of quantum hardware to get beyond the NISQ era, custom algorithms, and even generic hybrid solvers can become more competitive in the future.
+This may render quantum approaches important or even dominant for more complex disruptions and on larger scale.
+Figure 3: Synthetic experiments $line1$, $line2$, $line3$.
+Comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM.
 All the displayed solutions are feasible.
@@
 Panel a) - $line1$, results of both solvers were similar in terms of the objective, but the CPLEX computation was faster.
-Panel b) - $line2$ namely the double-track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective).
-On this example, there are instances, where the current CQM D-Wave hybrid solver outperforms CPLEX in terms of computational time.
-Increasing the value of the t_min parameter in most cases improves the quality of the solution at the cost of computational time.
+Panel b) - $line2$ namely the double-track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and the quantum solver (in terms of objective).
+In this example, there are instances where the current CQM D-Wave hybrid solver provides feasible, though not optimal solutions of similar quality to those that CPLEX provides when its computational time is limited to be similar to that of CQM.
+When CPLEX is required to solve the problem to optimality, it can be slower than CQM.
+Increasing the value of the t_min parameter does, in most cases, improve the quality of the solution at the cost of computational time.
 Panel c) - $line3$ all the displayed solutions are feasible.
-Results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
+The results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
 6.3 Real-life experiments on the considered rail network
@@
 We will also focus on the track closure situation, turning a double-track line segment into a single-track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2.
-Our computations address various use cases of train delays outside the network and within the network as well as the partial track closure.
+Our computations address various use cases of train delays outside the network and within the network as well as the track closures.
 We constructed $9$ networks with increased levels of difficulty ranging from small delays to disruptions, i.e.
 it is expected more trains are involved and/or disturbances are spread more over the network.
-These networks are tabulated in Table  2.
+Delays were assumed a priority fixed for each network manually according to everyday practice.
+These networks are tabulated in Table  6.
 Their details are as follows:
@@
   2. 2.
-    Networks with number $4$ or higher concern also disturbances due to closures.
+    Networks with number $4$ or higher also include disturbances due to closures.
     1. (a)
-      Networks $4$ and $5$ concerns rerouting trains from double-track line KTC-CB to the single-track line with higher passing times, see Fig.
-2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
+      Networks $4$ and $5$ concern rerouting trains from the double-track line KTC-CB to a single-track line with higher passing times, see Figure 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
     2. (b)
@@
 change of multiple line KZ-KO(STM) and double-track line KO-KL to single-track lines.
-(See Fig.
-2; the reason can be e.g.
-upcoming reconstruction works on this part of network.)
+(See Figure 2; the reason can be e.g.
+upcoming reconstruction works on this part of the network.)
     3. (c)
-      Networks $7$$9$ concerns closures of both $5$ and $6$, but each with different initial delays of trains.
-The intention of the last $3$ networks is to create a really challenging rescheduling problem.
+      Networks $7$ $9$ concern closures of both $5$ and $6$, but each with different initial delays of trains.
+Our intention with the last $3$ networks is to create a truly challenging rescheduling problem.
   3. 3.
@@
 1 10 KO-Ty IC 14006 KS 94113 Delayed IC (higher priority) is in conflict with on-time KS
-2 10 arr. Ty KS 94766 KS 40518 IC 41004 KS 44862 IC 4120 Delayed 5 trains on the arrival to Ty, random delays of several minutes
+2 10 arr. Ty KS 94766 KS 40518 IC 41004 KS 44862 IC 4120 Delayed 5 trains on the arrival to Ty, arbitrary delays of several minutes
 B. Disturbances originated within the network
-3 10 whole network 14 trains KS + IC Random delays of 14 trains delays up to 35 minutes
+3 10 whole network 14 trains KS + IC Arbitrary delays of 14 trains delays up to 35 minutes
 C. Disturbances within the network and closures
@@
 7 11 whole network all trains multiple-track KZ - KO(STM) and double-track KO - KL changed into single-track + double-track line KTC-CB closed as in 4 + delays as in 3
-8 11 whole network all trains Closures as in 7 + random delays of 13 trains up to 30 minutes
-9 11 whole network all trains Closures as in 7 + random delays of 15 trains up to 30 minutes
+8 11 whole network all trains Closures as in 7 + arbitrary delays of 13 trains up to 30 minutes
+9 11 whole network all trains Closures as in 7 + arbitrary delays of 15 trains up to 30 minutes
 Table 6: Real-life experiments.
 Networks of particular dispatching problems.
-In networks 00 - $5$ most of analyzed network is double-tracked, from network $6$ upward the number of single-track lines in the network increases at the expense of the number of double and multiple-track lines.
-For each network we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains.
-Beside the #$\#S$ stations there are $2$ depots in each network.
+In networks 00 - $5$ most of the analyzed network is double-tracked, from network $6$ upward the number of single-track lines in the network increases at the expense of double and multiple-track lines.
+In each experiment instance, initial delays were chosen manually, as arbitrary values plausible in practice.
+For each network, we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains.
+Besides the #$\#S$ stations there are $2$ depots in each network.
 n. CPLEX
@@
 # constr.
-obj. $\times$
-dmax
+obj.
 comp.
 time [s]
-obj. $\times$
-dmax
+obj.
 comp.
@@
 Observe that CQM always returns a feasible solution.
-For the networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX.
+For networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX.
 As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$).
-Meanwhile, those of CQM remain almost constant around $5.1s$.
+Meanwhile, those of CQM remain almost constant at around $5.1s$.
 QPU times remain not greater than $0.032s$.
-Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver even outperforms CPLEX with respect to computational time.
-It does not find the actual optimum, just a feasible solution close to it.
-For example, in network $9$, the computed CQM solution is around 30% worse than the optimal one.
-In conclusion, we expect benefits from the application of the hybrid solver on medium-scale railway networks with multiple closures, as in networks 7-9.
-This finding follows our observation on synthetic data in Fig.
-3 b), where the benefits appeared for more complex instances, too.
-To evaluate the extent of difficulty of the problems analyzed in Tab.
-3, Tab.
-4 presents mean (over stations and realizations) delays, resulting from CQM solutions, on selected stations (i.e.
-stations in the central part of the network in Fig.
-2).
-Meaningful delays are introduced starting from network $4$ onward, i.e.
-for networks with closures in a limited number of stations, see also Tab.
-2.
-From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Tab.
-2.
+Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver provides feasible solutions faster than exact CPLEX.
+It does not find the actual optimum though, just a feasible solution close to it.
+Certainly, there can be other (meta)heuristics that provide solutions of similar quality in an even shorter time, for instance, CPLEX itself when running for a limited time, as we have found in the case of the synthetic examples in Section 6.2.
+(This experience is also iniline with cases of other similar scheduling problems, see the work of [33] for another example.) Nevertheless, the experience with these examples suggests that CQM provides useful results in reasonable time, so it is at least one of the possible heuristics worth considering.
+We can expect benefits from the application of the hybrid solver on medium-scale railway networks with multiple closures, as in networks 7-9.
+This finding follows our observation on synthetic data in Figure 3 b), where the benefits appeared for more complex instances, too.
+To evaluate the extent of difficulty of the problems analyzed in Table 7, Table 8 presents the values of the delays averaged over trains and realizations, resulting from CQM solutions, for selected stations (i.e.
+stations in the central part of the network in Figure 2).
+Delays with a broader impact are introduced starting from network $4$ onward, i.e.
+for networks with closures in a limited number of stations, see also Table 6.
+From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Table 6.
 However, for networks $7$ $9$ the mean delays are the highest and almost uniformly distributed among all stations, showing that the spread of disruption impacts the whole network.
-These have appeared to be more challenging network disruptions.
-Fig.
-4 shows a more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (left, red) vs CQM (right, blue).
-For network $4$, bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller.
-This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Tab.
-2), and thus the spread is limited to these 3 stations.
+These appeared to be more challenging network disruptions.
+Figure 4 shows a more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (left, red) with exact CQM (right, blue).
+For network $4$, bigger delays occur mainly at stations KO(STM), KO, and CB for both CPLEX and CQM solutions.
+This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Table 6), and thus the spread is limited to these 3 stations.
 For network $7$, delays are more spread through the network due to disruption and initial disturbances.
-To conclude, we can expect an advantage from a hybrid (quantum-classical) approach, particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network.
+To conclude, we can expect an advantage from a hybrid (quantum-classical) approach, particularly in the case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network.
 To facilitate further evaluation of the solutions, train diagrams are presented in Appendix A.
@@
 Table 8: Real-life experiments.
-Mean (over trains and realizations) secondary delay resulted from CQM solutions (Tab.
-3) for all networks in Tab.
-2.
-Non-zero values for network 00 come from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes, delay but makes it up and leaves the network on time - these are typical issues for the railway traffic on analyzed network.
+Average (over trains and realizations) secondary delay resulted from CQM solutions (Table 7) for all networks in Table 6.
+Non-zero values for network 00 come from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes, delay but makes it up and leaves the network on time - these are typical issues for the railway traffic on the analyzed part of the network.
 Figure 4: Real-life experiments.
-Statistics of CPLEX (left, read) and CQM (right, blue) solutions (Tab.
-3) for selected networks in Tab.
-2.
+Statistics of CPLEX (left, read) and CQM (right, blue) solutions (Table 7) for selected networks in Table 6.
 The statistics were calculated over trains in the classical cases, and over trains and realizations in the CQM cases.
@@
 Let us analyze in more detail the role of the t_min solver parameter for selected problems.
-In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig.
-5 a) and b).
+In particular, for networks $7$ and $9$, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Figure 5 a) and b).
 The objective value is roughly log-linear in computational time.
 For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially.
-Overall, the improvement in the objective is overwhelmed by the increase of computational time.
+Overall, the improvement in the objective is overwhelmed by the increase in computational time.
 This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling.
-Such behavior is plausible in the case of QA methods (Soriani et al., 2022; Albash et al., 2017; Domino et al., 2022a).
+Such behavior is plausible in the case of QA methods [34, 1, 16].
 Figure 5: Real-life experiments.
-The sweep over t_min parameter for network $7$ a) and network $9$ b), see Tab.
-2.
+Dependence on the t_min parameter for network $7$ a) and network $9$ b), see Table 6.
 All solutions were feasible.
 For higher t_min we expect smaller objective values but higher total computational time.
-For each parameter value, there were $5$ realizations of the experiment were performed.
+For each parameter value, there were $5$ realizations of the experiment performed.
 Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
@@
 The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems.
-While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better.
+While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were competitive.
 This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
 While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models.
-In each computation, the CQM hybrid solver always has taken some advantage of the QPU.
-Hence, such a little help from the QPU must have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
+In each computation, the CQM hybrid solver always has actually used the QPU.
+Hence, such a little help from the QPU may have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
 Several future research directions can be determined.
-The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices (DiVincenzo, 2000) with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular).
-The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of quantum device to our advantage e.g.
+The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices [13] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular).
+The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of the quantum device to our advantage e.g.
 by modeling some stochastic behavior on railways.
-The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation (Goto, 2019).
+The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation [20].
 With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.
@@
 BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK).
-This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD).
-For the purpose of Open Access, the authors has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
+This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/02011 (LB) and 2023/07/X/ST6/00396 (KD).
+For the purpose of Open Access, the authors have applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
 We would like to thank Sebastian Deffner as well as the whole scientific community of the Department of Physics, University of Maryland, Baltimore County (UMBC) for valuable discussion; and to Bartłomiej Gardas, and Zbigniew Puchała for valuable motivation; and Katarzyna Gawlak, Akash Kundu, and Özlem Salehi for data validation and assistance with coding.
@@
 Appendix A. Train diagrams of CPLEX and CQM
-To further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX solution and two different quantum realizations; these can be seen in Fig.
-6 (left CPLEX, middle and right two realizations of CQM).
+To further evaluate the solutions, we present time-distance diagrams for network $7$, for the CPLEX optimal solution and two different quantum realizations; these can be seen in Figure 6 (left: exact CPLEX, middle and right: two realizations from CQM).
 We have chosen the line segment between GLC and KZ, offering a good demonstration.
-(We have introduced a small artificial ’distance’ between KO and KO(STM) which are contiguous locations for sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution.
-All the three proposed solutions have similar figures of merit, they all provide a valid option to be realized.
+(We have introduced a small artificial ’distance’ between KO and KO(STM) which are contiguous locations for the sake of better visibility, hence the horizontal lines in the diagram.) The red dotted lines represent the conflicted situation, and the green solid lines represent the solution.
+All three proposed solutions have similar figures of merit, they all provide a valid option to be realized.
 The conflicts are resolved in somewhat different ways, see e.g.
@@
   • in the CQM solution, second realization, 26103 passes first KTC-CB closure, then 5312 followed by 40150 go in opposite directions.
-This results in the following objective ×$\times d_{\max}$ values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$.
+This results in the following objective values: for CPLEX - $188.75$, for CQM first realization - $279$ for CQM second realization - $261.5$.
 Although the first solution is optimal, it is reasonable to supply dispatchers with the range of feasible solutions.
-Hence, these diagrams proved that the CQM hybrid solver can produce usable solutions for railway traffic management.
-An additional benefit of the probabilistic nature of the quantum-based method is that it produces different alternatives after each realization run without any modification.
+These diagrams illustrate that the CQM hybrid solver can produce usable solutions for railway traffic management.
+An additional benefit of the probabilistic nature of the quantum-based method is that it produces different alternatives after each sampling, without any modification.
 These could be offered to the dispatcher in a decision support system as possibilities.
@@
 Figure 6: Real-life experiment.
-Time-distance diagram solution of network $7$ CPLEX (left) and two realizations of CQM (middle and right), see Tab.
-2.
+Time-distance diagram solution of network $7$ CPLEX (left) and two realizations of CQM (middle and right), see Table 6.
 Real-life experiment.
-Time-distance diagram CQM solution of network $7$ (see Tab.
-2) first realization.
+Time-distance diagram CQM solution of network $7$ (see Table 6) first realization.
 Closures (yielding single-track line) are between CB and KTC as well as between KO(STM) and KZ.

### v6

@@
 Abstract
-We address the applicability of hybrid quantum-classical heuristics for practical railway rescheduling management problems.
-We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver as well as with CPLEX for comparison.
+We address the applicability of a hybrid quantum-classical heuristics for practical railway rescheduling management problems.
+We build an integer linear programming model and solve it with D-Wave’s quantum-classical hybrid solver (CQM) as well as with CPLEX, for comparison.
 The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments.
-All the requirements posed by the operator of the network are modeled.
-The computational results demonstrate the readiness for application and benefits of quantum-classical hybrid solvers in the realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation.
+All the requirements posed by the operator of the network included in the model.
+The computational results demonstrate the readiness for application and the benefits of quantum-classical hybrid solvers in a realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation.
 In particular, CQM as a probabilistic heuristic solver provides a number of feasible, close-to-optimal solutions the dispatcher can choose from.
@@
 Rail transport is expected to experience an increase in capacity demands due to changes in mobility needs resulting from climate policies, leading to traffic challenges in passenger and cargo rail transport.
-The situation is aggravated by the fact that the rolling stock and especially railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems.
-Rail transport, due to its technical and organizational characteristics is very sensitive to disturbances and disruptions in traffic: these extraordinary events have an impact on railway operations, typically resulting in delays [18].
+The situation is aggravated by the fact that the rolling stock and, especially, railway infrastructure cannot keep up with the increase in transport needs, which overloads railway systems.
+Rail transport, due to its technical and organizational characteristics, is very sensitive to disturbances and disruptions in traffic.
+These extraordinary events have an impact on railway operations, typically resulting in delays [18].
 Examples of such disturbances include late train departures/arrivals, extended dwell times, etc.
@@
 A disturbance is a smaller perturbation that can be handled solely by modifying the existing train paths, whereas a disruption affects rolling stock and crew schedules, too [5].
-In the present work, we focus on train rescheduling, which is the network operator’s perspective.
-Even though we also consider (partial) track closures, which are typically considered as disruptions, in our examples they will be managed by modifying existing train paths.
+In the present work, we focus on train rescheduling from the network operator’s perspective.
+Even though we also consider (partial) track closures, which are typically viewed as disruptions, in our examples they will be managed by modifying existing train paths.
 The impact of disturbances and disruptions can propagate to multiple sections in the railway network (c.f.
 [37] and references therein).
-Thus, ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers, and train operators.
-As a railway network is a complex non-local structure, modeling a bigger portion of it is necessary for the efficient suppression of the consequences of disturbances.
-While doing so, diverse objective functions were addressed, such as the (weighted) sum of delays [27], the maximal delay that cannot be avoided [10], or fuel consumption measures  [22].
-Large-scale rescheduling problems have to be addressed, and as the time available for making decisions is limited, they have to be solved almost in real time.
+Ensuring stable railway traffic and providing reliable service for passengers, rail cargo companies, and their clients is in the best common interest of railway infrastructure managers and train operators.
+A railway network is a complex non-local structure.
+Hence, modeling a bigger portion of it is necessary for the efficient suppression of the consequences caused by disturbances.
+Diverse objective functions have been employed for that purpose, such as the (weighted) sum of delays [27], the maximal delay that cannot be avoided [10], or fuel consumption measures  [22].
+These large-scale rescheduling problems have to be solved almost in real time, as the time available for making decisions is limited.
 The railway dispatching/rescheduling problem is recognized as equivalent to job-shop scheduling with blocking and no-wait constraints [36, 29].
-It has an extensive literature, already summarized in a number of reviews, including those by [5], by [26], or by [31].
-The maybe most frequently applied modeling strategy is based on alternative graphs [10, 29], resulting in a disjunctive program.
+This scheduling problem has an extensive literature, already summarized in a number of reviews, including those by [5], by [26], or by [31].
+One of the most frequently applied modeling strategies is based on alternative graphs [10, 29], resulting in a disjunctive program.
 Such a model can also be converted into a mixed integer linear program referred to as the big-M model.
-As this model often becomes too big to be solved on time, a number of strategies are applied, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and combinations of these; see the work of [26] for a review of the model and the methods.
-An alternative approach is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants.
+As this model often becomes too big to be solved on time, a number of strategies are applied in order to quickly find a suitable solution, including heuristics, specific branch-and-bound methods, alternative formulations, decompositions, and combinations of these; see the work of [26] for a review of the model and the methods.
+An alternative approach to this scheduling problem is to use time-indexed models: discrete-time units and binary decision variables that assign events to particular time instants.
 Even though the so-arising $0-1$ programs can be large, this method has been applied for both timetabling and dispatching/rescheduling [6, 28, 30].
 Meanwhile, $0-1$ programs are suitable as input for certain quantum hardware types.
-Indeed, job-shop problems addressed on hardware quantum annelaers [39], and even in certain hybrid approaches [25] use similar models, and our previous work [15] which was the first to apply a quantum solver to railway dispatching problems also followed this route.
+Indeed, job-shop problems addressed on hardware quantum annelaers [39], and even in certain hybrid approaches [25] use similar models.
+Our previous approach [15], which was the first to apply a quantum solver to railway dispatching problems, also followed this route.
 In the present work, however, we deal with a variant of the big-M model, albeit with discrete time variables, resulting in an integer linear program (ILP).
@@
 We do not include train cancellations in the model.
-We apply a new hybrid rescheduling algorithm combining classical-quantum modeling and based on Quantum Annealing (QA).
+We apply a new hybrid quantum-classical rescheduling algorithm combining classical heuristics with Quantum Annealing (QA).
 This work extends on a particular linear modeling strategy, partly explored on a toy model by [17].
 Our (ILP) model, tailored for the particular practical railway situation, is solved with proprietary D-Wave solvers as well as with CPLEX for comparison.
-We recognize several important scientific gaps.
+We recognize several important scientific gaps which the present work aims to contribute to.
 First, existing optimization models do not scale well: they can become huge and inefficient for larger real-life instances.
@@
 Second, the QA models for the studied optimization problem that have been introduced so far had been designed for pure QA implementations and have been demonstrated only on simple network setups due to the size limitations of the currently available Noisy Intermediate Scale Quantum (NISQ) devices.
-Third, no hybrid QA-based models have been used for real-life railway optimization so far, nor in other schedule-based modes, like public transport or air traffic planning/rescheduling.
+Third, to our knowledge, no hybrid QA-based models have been used for real-life railway optimization so far, nor in other schedule-based modes, like public transport or air traffic planning/rescheduling.
 In this paper, we demonstrate the quantum readiness of medium-scale railway rescheduling models: we successfully apply quantum methods in the rescheduling problem of an urban railway network.
@@
 Quantum computing devices became available for practical computational purposes in the last few years.
-Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system well-known in physics [23, 3].
-The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables, which is an NP-hard problem in general.
+Quantum annealers are such devices: they implement two-level quantum systems (quantum bits) with tunable pairwise interactions: an Ising system, well-known in physics [23, 3].
+The expression of the energy of such a system is one of the many equivalent forms of quadratic unconstrained binary optimization problems (QUBOs): the extremum of a quadratic function of binary variables.
+This is an NP-hard problem in general.
 The class of QUBO problems is well studied and has a number of applications; we refer the reader to a recent monograph edited by [32].
-Linear binary problems can also transformed to QUBOs; linear and quadratic constraints can be taken into account with penalties [21].
-Using appropriate binary encoding of the variables, integer programming problems can also be treated.
-Hence, an efficient QUBO solver can be useful in many problems, including the integer linear programs arising in the present contribution.
-Expressing a QUBO as the energy of an Ising system one gets a configuration of pairwise couplings of spins and local fields applied to individual spins, which can be realized in tunable quantum hardware.
-According to the adiabatic theorem of quantum mechanics [2], tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, under certain conditions the optimal configuration can be read out.
-Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, but it can be efficient in certain problems.
+Linear and quadratic constrained binary problems can also be transformed to QUBOs: linear and quadratic constraints can be taken into account with penalties [21].
+Using appropriate binary encoding of the variables, integer programming problems can be treated in this way as well.
+Hence, an efficient QUBO solver may be useful in many problems, including the integer linear programs arising in the present contribution.
+By expressing a QUBO as the energy of an Ising system, one gets a configuration of pairwise couplings of spins and local fields applied to individual spins, which, in turn, can be realized in tunable quantum hardware.
+According to the adiabatic theorem of quantum mechanics [2], tuning the system’s configuration from a suitably chosen initial state to the one corresponding to the actual objective slowly enough, the optimal configuration can be read out under certain conditions.
+Unfortunately, not all QUBOs meet the required conditions, so even in an ideal setting the quantum annealer is not a universal solver of hard problems, nevertheless it can be efficient in a variety of problems.
 Quantum annealing can be also viewed as a process analogous to simulated annealing, in which the noise required to avoid staying in local optima has quantum (as opposed to thermal) origins  [11].
-In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided, hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs.
+In real physical hardware, like the D-Wave machine, thermal noise and imperfections cannot be avoided.
+Hence, the physical devices realize a combination of thermal and quantum annealing, which renders them a probabilistic heuristic solver for QUBOs.
 They output a number of configurations typically close to, and possibly including the optimal one.
-The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits and not all pairs are coupled.
-The fixed topology of qubit couplings means that the problem’s graph defined by the nonzero coupling has to be embedded as an induced subgraph of the topology of the system.
-This embedding is a hard problem itself [41].
+The state-of-the-art physical quantum annealers are smaller systems of a few hundred quantum bits, where not all pairs of these are coupled.
+The fixed topology of qubit couplings means that the problem’s graph (defined by the nonzero couplings) has to be embedded as an induced subgraph of the graph describing the topology of the system.
+This embedding is a computationally hard problem itself [41].
 It often requires to couple multiple physical qubits to represent a logical bit of the problem.
@@
 Meanwhile, quantum technology keeps on developing, systems of bigger size and better topology are regularly announced, they are more and more affordable, and there is a growing community around them.
-Currently, the applicability of QA technologies is being explored.
+Currently, the applicability of QA technologies is being explored in various fields.
 In the railway context, the first applications of QA can be found in train dispatching/rescheduling [17] and rolling stock planning [4].
@@
 A follow-up paper by [17] laid down the principles of a more general modeling approach to railway rescheduling in light of QA, introducing a suitable QUBO / HOBO (Higher-Order Binary Optimization) encoding of these problems.
-The present work continues this line of research, which is gaining increasing attention, e.g.
+The present work continues this line of research, which is gaining increasing attention, see e.g.
 the recent work of [40].
@@
 These changes typically result in additional delays of trains, giving rise to secondary delays depending on the dispatching decision depending on the modification of the timetable.
-We will minimize the function of these secondary delays.
+We will minimize a function of these secondary delays.
 First we determine the subset of decision stations: these are the nodes of the network (typically: stations) where the events that are directly affected by the decisions can take place.
@@
 The goal of shunting is to move the train from the passengers’ service track to the depot or vice versa.
-The depot is treated as the station, and consider it as a black box, without a detailed layout.
+The depot is treated as a station, and consider it as a black box, without a detailed layout.
 We treat shunting movements as a service train from the depot to the starting station of the service train, or from the terminating station to the depot.
 Rolling stock circulation conditions are prescribed to ensure the precedence between the service train and the actual train.
-Our model shares features of both microscopic and macroscopic models.
-At the decision stations, it is microscopic as it takes into account station technology, track occupancy, rolling stock circulation, and shunting.
-Meanwhile, from the point of view of the rest of the network, it is a macroscopic model.
+Our model is mainly macroscopic while it contains some microscopic features, including station technology, track occupancy, rolling stock circulation, and shunting.
+Concerning switching zones, as no adjusting routes are modeled, independent parts of switching zones can be considered, thus parallel entries and departures can be performed.
 4 Methods and Model
@@
 $\mathcal{J}^{2\,(\text{track})}_{s}$ train pairs possibly competing for the same track of $s$
-$\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}$ train pairs possibly competing for an interlocking area of $s$ upon departure
-$\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon the departure of $j$ from $s$ while $j^{\prime}$ arrives to $s$ from the direction of $s^{\prime}$.
-$\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival from $s^{\prime}$ under special circumstances described in the main text.
-$\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ train pairs possibly competing for an interlocking area of $s$ upon arrival, under special circumstances described in the main text.
+$\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon departure
+$\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon the departure of $j$ from $s$ while $j^{\prime}$ arrives to $s$ from the direction of $s^{\prime}$.
+$\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon arrival from $s^{\prime}$ under special circumstances described in the main text.
+$\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ train pairs possibly competing for the same switch in an interlocking area of $s$ upon arrival, under special circumstances described in the main text.
 $\mathcal{C}^{2}_{j}$ station pairs subsequent in the route of $j$
@@
   • The sets $\mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly occupy the same track on the station $s$, thereby competing for the same track.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their departure from station $s$.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$.
-For these train pairs, the respective constraints will ensure that they do not pass the same interlocking area at the same time upon departure.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$.
-The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same interlocking area at the same time upon arrival, which is also enforced by the interlocking system.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their departure from station $s$.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$ (given trains’ paths fixed, $s^{\prime}$ is uniquely defined by $j^{\prime}$ and $s$).
+For these train pairs, the respective constraints will ensure that they do not pass the same switch in an interlocking area at the same time upon departure.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$.
+The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same switch in an interlocking area at the same time upon arrival, which is also enforced by the interlocking system.
 The possibility for M-O between the current and previous decision stations (e.g.
@@
   • The sets $\mathcal{C}^{2}_{j},\ \left(\forall j\in\mathcal{J}\right)$ are the sets of all station pairs that are subsequent in the route of $j$.
-  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction.
-  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions.
+  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction, and they cannot meet and overtake between the subsequent stations.
+  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs ($s$, $s^{\prime}$) that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions.
 The order within the pairs is determined by $j$.
+(Note that $s^{\prime}$ is uniquely defined by $j$, $j^{\prime}$ and $s$ because the routes of trains are fixed.)
 All these sets can be enumerated based on the input data in a straightforward manner. A brief summary of the sets is provided in Table 1.
@@
 Interlocking area occupancy
-These ensure that trains cannot meet in interlocking areas:
+These ensure that trains that use the same switch cannot meet in interlocking areas:
 $t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$
@@
 Let us discuss the number of variables and constraints in our model.
-The number of time variables #$\#t$, the product #$\#\mathcal{J}\#\mathcal{S}$ of the number of trains and the number of stations serves as a trivial but rough upper bound.
-(See also Section 3.1 of the work of  [17]).
-To obtain a more accurate estimate of the number of these variables we observe that in our model no train visits all the stations.
-To provide an estimate, we introduce $\alpha\in(0,1)$ as the fraction of the stations visited by an average train.
-Thus we have
-#$\#(t)\approx\alpha\#\mathcal{J}\#\mathcal{S}\leq\#\mathcal{J}\#\mathcal{S}.$ (20)
-as a tighter bound.
-In what follows we will mostly $\alpha=1$, the rough bound, however, we will keep the parameter $\alpha$ in the formulae to reflect its role in the scaling.
-Let us assume, that half the of trains (#$\frac{\#\mathcal{J}}{2}$) are going in one direction and the other half are going in the other.
-Then the number of trains each one meets going in the same direction is $\alpha\frac{\#\mathcal{S}\#\mathcal{J}}{2}$.
-This yields #$\#\mathcal{J}\cdot\frac{\alpha\#\mathcal{S}\#\mathcal{J}}{2}$ $y^{(\text{out})}$ variables.
-However, because of the symmetry property in Equation (19), only half of these variables are independent, thus we have $\alpha\#\mathcal{S}\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{out})}$ of these variables.
-As for the trains going in opposite directions, we assume that each train meets at most #$\frac{\#\mathcal{J}}{2}$ trains on each station in the single-track parts of the network.
-Hence, the number of $z$ variables is at most #$\#\mathcal{J}\cdot\frac{1}{2}\cdot\frac{\alpha\#\mathcal{J}(\#\mathcal{S}-1)}{2}$ (here, #$\#\mathcal{S}-1$ comes from the observation that the $z$-variables are tied to pairs of stations).
-Variables $y^{(\text{in})}$ occur mainly in Eqs.
-(4.2.2)-(4.2.2), relevant for a few stations only: there is typically one such station in the network, yielding at most #$\frac{\#\mathcal{J}^{2}}{4}$ $y^{(\text{in})}$ variables.
-In summary, the number of precedence variables can be estimated as
-#$\begin{split}&\#y^{(\text{out})}\leq\alpha\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\leq\frac{\#\mathcal{J}^{2}}{4}\#\mathcal{S}\\ &\#y^{(\text{in})}\leq\frac{\#\mathcal{J}^{2}}{4}\\ &\#z\leq\alpha\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\leq\frac{\#\mathcal{J}^{2}}{4}(\#\mathcal{S}-1)\\ &\#y+\#z=\#y^{(\text{out})}+\#y^{(\text{in})}+\#z<\frac{\#\mathcal{J}^{2}}{2}\#\mathcal{S}\end{split}$ (21)
+To assess the problem’s size we will use here a worst-case approximation, namely, we assume temporarily that all trains visit all stations and any pair of trains can meet each other; these will be assumed throughout the entire consideration on scaling.
+Under such assumptions the number of time variables #$\#t$ can be upper bounded by the product #$\#\mathcal{J}\#\mathcal{S}$ of the number of trains and the number of stations:
+#$\#(t)\leq\#\mathcal{J}\#\mathcal{S}.$ (20)
+Each train meets at most (#$(\#\mathcal{J}-1)\#\mathcal{S}$ trains going in same direction, which results in #$\#\mathcal{J}(\#\mathcal{J}-1)\#\mathcal{S}$ $y^{(\text{out})}$ variables in the model.
+However, because of the symmetry property in Equation (19), only half of these variables are independent, thus we have at most #$\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}$ of these variables.
+As for the trains going in opposite directions, each train meets #$\#\mathcal{J}-1$ trains at each station.
+(Although we have a quadruple of indices $(j,j^{\prime},s,s^{\prime})$ in the definition of $z$ variable, as routes of trains are fixed within each model instance, the elements of this quadruple are not independent.
+In particular, the index $s^{\prime}$ is uniquely defined by the other indices).
+Hence, the number of $z$ variables is at most #$\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}$ (having taken into account the symmetry property in Equation (19) again).
+Variables $y^{(\text{in})}$ occur mainly in Equations (4.2.2)-(4.2.2), and are relevant relevant for some stations only.
+Nevertheless, in the worst case there can be one such variable assigned for each pair of trains and each station, yielding #$\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}$ $y^{(\text{in})}$ variables after the reduction by symmetry properties.
+In summary, the upper bound for the number of precedence variables reads
+#$\begin{split}\#y^{(\text{out})}&\leq\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}\\ \#y^{(\text{in})}&\leq\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}\\ \#z&\leq\frac{\#\mathcal{J}(\#\mathcal{J}-1)\cdot\#\mathcal{S}}{2}\\ \#y+\#z&=\#y^{(\text{out})}+\#y^{(\text{in})}+\#z\leq\frac{3}{2}\#\mathcal{J}(\#\mathcal{J}-1)\#\mathcal{S}\end{split}$ (21)
 Note that these bounds are rather pessimistic: setting the parameter $d_{\max}$ as an upper bound on the secondary delays implies a reduction in the number of trains that can potentially conflict.
 However, the actual number of them can depend on the train density and structure of the timetable in various ways, hence, it is not possible to provide a tighter bound that would be valid in general.
-Finally, let us remark as the $z$-variables are mainly tied to single-track traffic, the model is more complex for networks dominated by single-track segments.
-Based on Eqs. (20) and Equation (21), the scaling of the upper bound on the the total number of decision variables can be estimated by
-#$\#vars=\#t+\#y+\#z\approx\text{const}\#\mathcal{J}^{2}\#\mathcal{S}+\text{const'}\#\mathcal{J}\#\mathcal{S},$ (22)
+Finally, let us remark that as the $z$-variables are mainly tied to single-track traffic, the model is more complex for networks dominated by single-track segments.
+Based on Equations (20) and (21), the scaling of the upper bound on the total number of decision variables can be estimated by
+#$\#vars=\#t+\#y+\#z\approx O(\#\mathcal{J}^{2}\#\mathcal{S}),$ (22)
 which is quadratic in the number of trains and linear in number of stations.
 Now let us estimate the number of constraints.
-Considering the headway constraints in Equation (4), there is one such constraint for each variable.
-However, because of the symmetry property in Equation (19), only half of the variables are independent, hence, the number of these constraints is twice the number of independent variables.
+Considering the headway constraints in Equation (4), there are two such constraints for each independent $y^{(\text{out})}$ variable.
+(Because of the symmetry property in Equation (19), only half of the $y^{(\text{out})}$ variables are independent).
 Along the same arguments, in case of the constraints in Equation (11), i.e.
-station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each $z$ variable.
-It is more complicated to estimate the number of interlocking area occupation constraints in Eqs.
-(4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables.
-A good assumption is to consider $2$ constraints for each $y^{(\text{out})}$ variable, $2$ constraints for each $y^{(\text{in})}$ variable, and $2$ constraints for each $z$ variable.
-Further, we expect one minimal dwell time constraint per train and station (described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term).
-Concerning train order constraints in Eqs (17) (18), we also expect one constraint per train and station (also described by a $\alpha\#\mathcal{J}\#\mathcal{S}$ term).
-We exclude the rolling stock circulation constraints from this consideration as their contribution to the number of constraints is very limited.
-In summary, the total number of constraints obeys the following upper bound:
-#$\#constr.<6\#y^{(\text{out})}+2\#y^{(\text{in})}+4\#z+2\alpha\#\mathcal{J}\#\mathcal{S}.$ (23)
-Substituting the estimated number of variables in Equation (21) we have
-#$\#constr.<\alpha\frac{3}{2}\#\mathcal{J}^{2}\#\mathcal{S}+\frac{1}{2}\#\mathcal{J}^{2}+\alpha\#\mathcal{J}^{2}(\#\mathcal{S}-1)+2\alpha\#\mathcal{J}\#\mathcal{S}$ (24)
-Setting $\alpha=1$ as the worst case, we get
-#$\#constr.<\frac{5}{2}\#\mathcal{J}^{2}\#\mathcal{S}-\frac{1}{2}\#\mathcal{J}^{2}+2\#\mathcal{J}\#\mathcal{S}.$ (25)
+station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each independent $z$ variable.
+It is more complicated to estimate the number of interlocking area occupation constraints in Equations (4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables.
+A good assumption is to consider $2$ constraints for each independent $y^{(\text{out})}$ variable, $2$ constraints for each independent $y^{(\text{in})}$ variable, and $2$ constraints for each independent $z$ variable.
+Further, we expect one minimal dwell time constraint per train and station, resulting in #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case.
+Concerning train order constraints in Equations (17) and (18), there can be one constraint per train and station, which yields additional #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case.
+As we do not consider splitting and joining trains on the route, at most one rolling stock circulation constraint arises per train, yielding at most #$\#\mathcal{J}$ such constraints.
+In summary, the total number of constraints obeys the following (worst-case) upper bound:
+#$\#constr.\leq 6\#y^{(\text{out})}+2\#y^{(\text{in})}+4\#z+2\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (23)
+Substituting the estimated number of variables in Equation (21) we have:
+#$\#constr.\leq 6\#\mathcal{J}^{2}\#\mathcal{S}-4\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (24)
 Let us stress that the provided generic pessimistic upper bounds only serve the purpose of getting an overall picture of the scaling; the particular number of variables and constraints is lower and will be specified for our problem instances.
@@
 upper lim
-1 3 59 118 180 1538 5221 7889 24721
-2 5 39 142 200 1394 3802 8491 18642
-3 5 21 79 110 592 1102 3057 5502
+1 3 59 118 177 1538 15399 7889 62009
+2 5 39 142 195 1394 11115 8491 44889
+3 5 21 79 105 592 3150 3057 12831
 Table 5: Estimated and actual problem sizes on $line1$ $line2$ and $line3$ scenarios. For each scenario, the mean over instances was computed.
@@
 While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were competitive.
-This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
+These results support the expectation that future quantum devices can become efficient in solving large-scale problems, like, in our case, real-time railway dispatching problems on the national scale, even in the range that is beyond the scope of current exact models and heuristics.
 While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models.
-In each computation, the CQM hybrid solver always has actually used the QPU.
-Hence, such a little help from the QPU may have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
-Several future research directions can be determined.
-The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices [13] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular).
+It is important to note that the CQM hybrid solver has always used the QPU in each computation.
+The actual assistance from the QPU may have boosted the classical heuristics in the hybrid solver.
+A possible way of improving on the presented technique would be to find a way of determining the smallest suitable $d_{\text{max}}$ based on the input data instead of an ad-hoc choice, by using some fast enough methods.
+Apart from that, several future research directions can be determined.
+The first concerns the ongoing worldwide efforts on the development of fault-tolerant quantum computing devices [13] with proper error correction/mitigation.
+Those future quantum devices are expected to boost the efficiency of approaches supported by quantum computing (and QA in particular).
 The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of the quantum device to our advantage e.g.
 by modeling some stochastic behavior on railways.
-The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation [20].
+The third concerns the evolution and application of quantum-inspired techniques such as tensor network techniques or simulated bifurcation [20].
 With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.

### v7

@@
 The proposed approach is demonstrated on a real-life heterogeneous urban network in Poland, including both single- and multi-track segments.
-All the requirements posed by the operator of the network included in the model.
+All the requirements posed by the operator of the network are included in the model.
 The computational results demonstrate the readiness for application and the benefits of quantum-classical hybrid solvers in a realistic railway scenario: they yield acceptable solutions on time, which is a critical requirement in a rescheduling situation.
@@
 Rolling stock circulation conditions are prescribed to ensure the precedence between the service train and the actual train.
-Our model is mainly macroscopic while it contains some microscopic features, including station technology, track occupancy, rolling stock circulation, and shunting.
-Concerning switching zones, as no adjusting routes are modeled, independent parts of switching zones can be considered, thus parallel entries and departures can be performed.
 4 Methods and Model
@@
 4.1.1 Sets
-$\hat{S}$ stations
-$S$ decision stations
-$J$ trains
+$\hat{\mathcal{S}}$ stations
+$\mathcal{S}$ decision stations
+$\mathcal{J}$ trains
 $\mathcal{S}_{j}$ stops of train $j$ (ordered)
@@
 Table 1: Summary of sets
-Let $\hat{S}$ denote the set of all stations, and $S\subset\hat{S}$ the decision stations which have been selected in advance, as already described.
+Let $\hat{\mathcal{S}}$ denote the set of all stations, and $\mathcal{S}\subset\hat{\mathcal{S}}$ the decision stations which have been selected in advance, as already described.
 The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path.
@@
   • The sets $\mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall s\in\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their departure from station $s$.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$ (given trains’ paths fixed, $s^{\prime}$ is uniquely defined by $j^{\prime}$ and $s$).
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ in order to have $j$ to depart from $s$ while $j^{\prime}$ arrive at $s$ from the direction of $s^{\prime}$ (given trains’ paths fixed, $s^{\prime}$ is uniquely defined by $j^{\prime}$ and $s$).
 For these train pairs, the respective constraints will ensure that they do not pass the same switch in an interlocking area at the same time upon departure.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
-  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-O possibility for them between $s$ and $s^{\prime}$.
+  • The sets $\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})},\ \left(\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}\right)$ are the sets of train pairs that can possibly compete for the same switch in an interlocking area of $s$ upon their arrival at $s$ so that either both of them come from the direction of $s^{\prime}$ but there is a M-O possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$.
 The constraints related to these sets and those related to $\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}}$ together will ensure that pairs of inbound trains do not use the same switch in an interlocking area at the same time upon arrival, which is also enforced by the interlocking system.
@@
   • The sets $\mathcal{C}^{2}_{j},\ \left(\forall j\in\mathcal{J}\right)$ are the sets of all station pairs that are subsequent in the route of $j$.
-  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction, and they cannot meet and overtake between the subsequent stations.
-  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)$ are those of all station pairs ($s$, $s^{\prime}$) that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions.
+  • The set $\mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}\times\mathcal{J}\right)$ are the sets of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction, and they cannot meet and overtake between the subsequent stations.
+  • The sets $\mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}},\ \left(\forall(j,j^{\prime})\in\mathcal{J}\times\mathcal{J}\right)$ are those of all station pairs ($s$, $s^{\prime}$) that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions.
 The order within the pairs is determined by $j$.
@@
 $f(t)=\sum_{j\in\mathcal{J}}w_{j}\left(t^{(\text{out})}(j,s^{*})-\upsilon(j,s^{*})\right),$ (2)
-where $s^{*}$ is the last element of $S_{j}$, and $w_{j}$ are weights for each train representing its priority.
+where $s^{*}$ is the last element of $\mathcal{S}_{j}$, and $w_{j}$ are weights for each train representing its priority.
 The weights have to be chosen in such a way that they reflect the rules and practices of the local dispatching system.
@@
 Each train has to occupy the station node for a prescribed time duration at each station:
-$t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s),\ \forall j\in\mathcal{J},\ \forall s\in S_{j}.$ (10)
+$t^{(\text{out})}(j,s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{dwell})}(j,s),\ \forall j\in\mathcal{J},\ \forall s\in\mathcal{S}_{j}.$ (10)
 Station track occupancy
 Station tracks can be occupied by at most one train at a time:
-$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\ {\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}},\$ (11)
+$t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\ {\forall s\in\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}},\$ (11)
 where $C$ is chosen similarly to that in Equation (4) again.
@@
 $t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),$
-$\forall s\in S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s};$ (12)
+$\forall s\in\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out})}_{s};$ (12)
 $t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot z(j^{\prime},j,s^{\prime},s),$
-$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}};$ (13)
+$\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}};$ (13)
 $t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{out})}(j^{\prime},j,s^{\prime}),$
-$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}};$ (14)
+$\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{noMO})}_{s,s^{\prime}};$ (14)
 $t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{(\text{switch})}(j,j^{\prime},s)-C\cdot y^{(\text{in})}(j^{\prime},j,s),$
-$\forall(s,s^{\prime})\in S^{\times 2},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}},$ (15)
+$\forall(s,s^{\prime})\in\mathcal{S}\times\mathcal{S},\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{in},\text{MO})}_{s,s^{\prime}},$ (15)
 with a choice of $C$ similar again to that in Equation (4). These constraints are also necessary for to make the realization of the modified timetable possible using the interlocking systems.
@@
 If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
-$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s),\ \forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}.$ (16)
+$t^{(\text{out})}(j^{\prime},s)\geq t^{(\text{in})}(j,s)+\tau^{\text{(turn)}}(j,j^{\prime},s),\ \forall s\in\mathcal{S}\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{turn})}_{s}.$ (16)
 Order of trains
@@
 Formally:
-$\begin{split}y^{(\text{in})}(j,j^{\prime},s^{\prime})=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(s,s^{\prime})\in S\times S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{\text{in}},\text{MO})}_{s^{\prime},s}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}.\end{split}$ (18)
+$\begin{split}y^{(\text{in})}(j,j^{\prime},s^{\prime})=y^{(\text{out})}(j,j^{\prime},s^{\prime}),\\ \forall(s,s^{\prime})\in\mathcal{S}\times S,\ \forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{switch},\text{\text{in}},\text{MO})}_{s^{\prime},s}\cap\mathcal{J}^{2\,(\text{track})}_{s^{\prime}}.\end{split}$ (18)
 The objective function in Equation (2), together with the constraints in Equation (3)-(18) define our mathematical programming model.
@@
 Now let us estimate the number of constraints.
-Considering the headway constraints in Equation (4), there are two such constraints for each independent $y^{(\text{out})}$ variable.
+Considering the headway constraints in Equation (4), there are at most two such constraints for each independent $y^{(\text{out})}$ variable.
 (Because of the symmetry property in Equation (19), only half of the $y^{(\text{out})}$ variables are independent).
 Along the same arguments, in case of the constraints in Equation (11), i.e.
-station track occupancies, there are typically $2$ of these for each independent $y^{(\text{out})}$ variable, and for the single-line track occupancy constraints in Equation (9) there are $2$ constraints per each independent $z$ variable.
-It is more complicated to estimate the number of interlocking area occupation constraints in Equations (4.2.2))-(4.2.2)) as they involve $z$, $y^{(\text{out})}$ and $y^{(\text{in})}$ variables.
-A good assumption is to consider $2$ constraints for each independent $y^{(\text{out})}$ variable, $2$ constraints for each independent $y^{(\text{in})}$ variable, and $2$ constraints for each independent $z$ variable.
+station track occupancies, there are at most $2$ of these for each independent $y^{(\text{out})}$ variable.
+Analogously, for the single-line track occupancy constraints in Equation (9) there are at most $2$ constraints per each independent $z$ variable.
+Furthermore, in the worst case there are: $2$ constraints for each independent $y^{(\text{out})}$ variable from Equation (4.2.2), $2$ constraints for each independent $y^{(\text{out})}$ variable from Equation (4.2.2), $2$ constraints for each independent $z$ variable from Equation (4.2.2), and $2$ constraints for each independent $y^{(\text{in})}$ variable from Equation (4.2.2).
 Further, we expect one minimal dwell time constraint per train and station, resulting in #$\#\mathcal{J}\#\mathcal{S}$ constraints in the worst case.
@@
 Substituting the estimated number of variables in Equation (21) we have:
-#$\#constr.\leq 6\#\mathcal{J}^{2}\#\mathcal{S}-4\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (24)
+#$\#constr.\leq 7\#\mathcal{J}^{2}\#\mathcal{S}-5\#\mathcal{J}\#\mathcal{S}+\#\mathcal{J}.$ (24)
 Let us stress that the provided generic pessimistic upper bounds only serve the purpose of getting an overall picture of the scaling; the particular number of variables and constraints is lower and will be specified for our problem instances.
@@
 upper lim
-1 3 59 118 177 1538 15399 7889 62009
-2 5 39 142 195 1394 11115 8491 44889
-3 5 21 79 105 592 3150 3057 12831
+1 3 59 118 177 1538 15399 7889 72275
+2 5 39 142 195 1394 11115 8491 52299
+3 5 21 79 105 592 3150 3057 14931
 Table 5: Estimated and actual problem sizes on $line1$ $line2$ and $line3$ scenarios. For each scenario, the mean over instances was computed.